### v2

# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

Abstract
Large language models (LLMs) frequently hallucinate due to misaligned self-awareness, generating erroneous outputs when addressing queries beyond their knowledge boundaries. While existing approaches mitigate hallucinations via uncertainty estimation or query rejection, they suffer from computational inefficiency or sacrificed helpfulness. To address these issues, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework, integrating fast and slow reasoning systems to harmonize reliability and usability. The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate use of high-confidence outputs. For uncertain predictions, a slow refinement model conducts targeted reasoning to improve accuracy. To align model behavior with our proposed object, we propose a hybrid training pipeline, enhancing self-awareness without degrading task performance. Evaluations on dialogue state tracking tasks demonstrate that EKBM achieves superior model reliability over uncertainty-based baselines. Further analysis reveals that refinement substantially boosts accuracy while maintaining low computational overhead. Our work establishes a scalable paradigm for advancing LLM reliability and balancing accuracy and practical utility in error-sensitive applications.
Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling
Hang Zheng1††thanks: Equal contribution., Hongshen Xu1, Yuncong Liu1, Lu Chen1, Pascale Fung2, Kai Yu1 1X-LANCE Lab, Department of Computer Science and Engineering, Shanghai Jiao Tong University 2Center for Artificial Intelligence Research (CAiRE), Hong Kong University of Science and Technology {azure123, xuhongshen, chenlusz, kai.yu}@sjtu.edu.cn
1 Introduction
Recently, large language models (LLMs) have exhibited impressive text generation capabilities Abdullah et al. (2022). However, despite their advancements, LLMs are prone to hallucination, where generated content misaligns with context or factual information Zhang et al. (2023). Such inaccuracies can be particularly harmful in low error-tolerance applications, eroding user trust in LLM reliability.
Extensive works have been devoted to mitigating hallucinations in LLMs Tonmoy et al. (2024). Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, leading to errors when models attempt to answer questions beyond their expertiseLi et al. (2024b). Improving the model’s self-awareness—its capacity to accurately assess its own knowledge boundaries and outputs correctness, can effectively mitigate this misalignment. Several uncertainty-based methods have been proposed to estimate model confidence and indirectly reflect the model’s self-awareness Huang et al. (2024). However, these approaches are not inherent to the model’s endogenous capabilities and frequently suffer from high computational costs or instability across thresholds. Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries Xu et al. (2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
To mitigate these limitations, we propose a novel alignment objective: improving model awareness for knowledge boundaries, enabling models to respond maximally while explicitly distinguishing outputs of high and low confidence. We introduce the concepts of "sure" and "unsure" to clarify this distinction. A reliable system should deliver near-perfect accuracy for "sure" predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing immediate utilization of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to enhance accuracy of low-confidence predictions. Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage. The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective operation.
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately categorize its outputs into "sure" and "unsure" categories compared to baseline methods. Additionally, we develop an automatic refinement model for "unsure" outputs, showing that the EKBM system achieve superior performance compared to traditional supervised fine-tuning (SFT) approaches. Our analysis of various training methods and model behaviors provides further valuable insights.
Our contributions are summarized as follows:
We present the EKBM framework, integrating fast-thinking confidence labeling with slow-thinking refinement, along with metrics for evaluating LLM self-awareness and reliability.
We develop a comprehensive training pipeline that enhances LLM self-awareness, yielding more reliable models.
We perform extensive experiments to demonstrate the effectiveness and scalability of the EKBM framework, providing valuable insights for future research.
2 Related Work
The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs Yin et al. (2024); Li et al. (2024b). Misalignment between model behavior and knowledge boundaries can result in factual hallucinations Huang et al. (2023) and ambiguous responses Liu et al. (2023).
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities Manakul et al. (2023) and consistency Chen and Mueller (2024). Calibration strategies, including prompting Tian et al. (2023) and fine-tuning Tao et al. (2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations Li et al. (2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes Huang et al. (2024); Zhou et al. (2024b).
To address this misalignment, researchers have implemented knowledge-enhanced fine-tuning Zhou et al. (2024a) and retrieval-augmented generation (RAG, Lewis et al., 2020) to bolster model capabilities, thereby reducing hallucinations. Some approaches also guide models in self-reflection to minimize self-contradictions Wei et al. (2022); Ji et al. (2023), albeit at the cost of increased overhead. For queries beyond the model’s capabilities, refusal or rejection strategies can prevent misinformation Xu et al. (2024b); Chen et al. (2024), though they may inadvertently reduce overall helpfulness.
3 Problem Formulation
3.1 Alignment for Reliability
To enhance the reliability of LLMs, a prevailing method involves aligning their performance with high-quality training data. This alignment aims to achieve the following objective:
where denotes a scoring function, is the input prompt, and and denote correct and incorrect responses, respectively, ensuring the score of correct responses exceeds that of incorrect ones.
Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries. For unanswerable queries, the model is encouraged to refrain from responding. The corresponding alignment objective can be articulated as:
where denotes a truthful rejection, and an incorrect response. This suggests that refusal is preferable to providing an incorrect answer.
While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios. Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence ("sure") and low-confidence ("unsure") responses. The model must ensure high accuracy in the "sure" category, while tolerating some errors in the "unsure" category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and represent correct and incorrect responses, respectively, while and represent "sure" and "unsure" as levels of confidence.
This objective guarantees that correct "sure" predictions are prioritized, followed by correct "unsure" predictions, which, although less confident, retain utility. Incorrect predictions in the "unsure" category are discouraged but more acceptable than those in the "sure" category, as the latter significantly undermine reliability.
3.2 Reliability Evaluation
In this work, we focus on complex multi-slot problems and have modified several metrics for better evaluation of model reliability.
3.2.1 Weighted-F1
F1-score is usually taken as the evaluation metric for multi-slot problems:
Where TP, FP and FN represent True Positives, False Positives and False Negatives respectively.
We modify the F1 score to adapt to the alignment objective of our EKBM framework. Since predictions labeled as "unsure" represent an intermediate state of uncertainty, we assign a weighted contribution (denoted by and ) to predictions in this category. The weights reflect the partial utility of "unsure" predictions as potential candidates for refinement. The modified precision and recall are defined as:
Where STP and SFP represent Sure True Positives and Sure False Positives, while UTP and UFP denote the corresponding cases of Unsure. For UTP in Recall calculation, we only require that a prediction is partially correct (able to be refined). The modified Weighted-F1 score is defined as:
The design allows flexible handling of "unsure" predictions, effectively managing their intermediate utility. influences the calculation of Precision, reflecting the model’s tolerance for errors among "unsure" predictions; as increases, tolerance for such errors decreases. Conversely, affects Recall, indicating the degree of reward for providing additional reference information for "unsure" predictions. As increases, the model is encouraged to enhance helpfulness by generating more "unsure" predictions as a basis for refinement.
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all "unsure" predictions. (2) = 0, = 1: Errors in the "unsure" category are ignored, rewarding effective recall, representing the theoretical upper limit post-perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
3.3 Reliability Metrics
We employ two specific Weighted-F1 values as performance metrics:
Optimal-F1 Defined as , representing the theoretical upper limit after perfect refinement of "unsure" predictions.
Quality-F1 Defined as , it assigns "unsure" predictions half the weight of "sure" ones, encouraging the model to select "sure" when confident while minimizing penalties for "unsure" outputs. Quality-F1 measures the model’s self-awareness in accurately classifying both types.
Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
4 Method
4.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the "unsure" predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages.
Fast Prediction with Confidence Labeling
In the first stage, the model performs the primary task while simultaneously assigning a confidence label ("sure" or "unsure") to each prediction. Predictions labeled as "sure" are directly accepted as final outputs, ensuring high precision.
Slow Refinement for Unsure Prediction
In the second stage, the "unsure" predictions are further refined. This refinement process can involve various strategies, such as user confirmation, automated multi-step reasoning, or other post-processing techniques. This additional stage incurs extra cost but increases the overall accuracy and reliability of the system.
Our two-stage framework balances immediate usability with comprehensive coverage, making the model both reliable and helpful.
4.2 Reliability Alignment
To align the model’s behavior with our proposed alignment objective, we explore various training approaches to enhance its intrinsic self-awareness Xu et al. (2025). To minimize interference from task performance improvements, we introduce minimal supervisory signals related to task knowledge. We employ Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), with the data construction process illustrated in Figure 2 (2).
SFT Data Construction
We assume that task-specific knowledge boundary of the model can be approximated by repeatedly sampling the model’s outputs on the training data. Based on this, our SFT data construction approach is intuitive. We conduct multiple rounds of sampling for each training sample. Predictions that are consistently accurate are labeled as "sure," while those exhibiting errors or omissions receive an "unsure" label. Importantly, we refrain from correcting erroneous outputs to avoid introducing supervisory signals that could compromise the model’s task performance.
The number of sampling rounds is a hyperparameter denoted as i. A larger i results in a more conservative model, as the proportion of "unsure" labels in the dataset increases with greater i. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
DPO Preference Data Construction
DPO enables the model to learn directly from preference pairs, making the design of an effective preference strategy crucial. Given our alignment objective, it is intuitive to use Weighted-F1 as the preference score function. The values in Weighted-F1 significantly influence the model’s behavior and performance. Through experimentation (see Section 5.6), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement. Therefore, this setting will be used by default unless specified. To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling, aligning with the model’s output distribution.
4.3 Refinement Model
We train automated refinement models for each dataset to further optimize by reasoning on unsure predictions, utilizing supervised fine-tuning (SFT) based on the LLAMA3 8B model. Overall, our refinement process operates at a fine-grained level, with each unsure prediction being individually refined. We employ a multi-step reasoning paradigm similar to the Chain of Thought Wei et al. (2022) methodology and use GPT-4o Hurst et al. (2024) to generate the reasoning process. This process is illustrated in Figure 2 (3), and detailed information can be found in Appendix A.
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Dubey et al. (2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B Yang et al. (2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
5 Experiments
We conduct experiments focusing on three core research questions:
RQ1 Does our proposed reliability training pipeline outperform current methods in enhancing self-awareness abilities?
RQ2 Does our EKBM framework improve the overall performance of LLMs on complex tasks?
RQ3 Does our approach demonstrate strong scalability, performing well across foundation models with various performance?
5.1 Experiments Setup
5.1.1 Model and Baselines
For our experiments, we utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone.
For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO.
To evaluate the reliability of our proposed methods, we compare their performance against several baseline approaches, including three prompt-based and three uncertainty-based methods.
Prompt-based
methods include three approaches: Direct, Verbose, and Self-Reflection (SR). The Direct method predicts outputs without incorporating any confidence labels, while Verbose generates confidence labels alongside predictions based on predefined principles. SR evaluates prediction reliability through individual reflections.
Uncertainty-based
methods consist of three approaches: Token Probability (Prob), Self-Consistency (SC), and P(True) Kadavath et al. (2022). These methods classify outputs as "sure" or "unsure" using different techniques. Prob calculates average token-level probabilities, SC assesses the frequency of predictions across multiple samples, and P(True) determines the probability of the "True" token during model self-evaluation.
The details of the baselines can be found in E.1.
5.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology Xu et al. (2024a). Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.
To ensure a robust evaluation, we select three datasets: MultiWOZ-2.4 Ye et al. (2021), BiTOD Lin et al. , and SGD Rastogi et al. (2020). For BiTOD we only utilize the English version. For SGD, we randomly select 10,000 test samples.
5.1.3 Training Details
We modify the DeepSpeed-Chat Yao et al. (2023) framework to accommodate our experimental setup. We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others. By default, we utilize 1,000samples for Reliability-SFT Training and 2,000samples for DPO Training. All models are trained for 1 epoch using NVIDIA A800 GPUs.
5.2 Reliability Evaluation
The results of our experiments on model reliability are summarized in Table 1. We compare our reliability-training pipeline with baseline algorithms, emphasizing the model’s self-awareness ability. For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, markedly enhancing self-awareness.
Our approach consistently achieves higher Quality-F1 scores, indicating improved effectiveness in categorizing predictions as "sure" or "unsure" based on confidence. Notably, a high Precisionsure indicates reliable "sure" predictions, fostering user trust with minimal verification. Simultaneously, a high Recalltotal ensures that the "unsure" category encompasses a broad range of potentially correct outputs, forming a solid basis for further refinement. Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In other words, our system offers improved performance by effectively balancing precision and recall, ensuring both immediate reliability and the potential for improved post-refinement accuracy.
In our approach, methods incorporating DPO generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores. This suggests that Joint Training models are more conservative, relying on subsequent refinement, whereas Post Training models exhibit greater confidence, yielding predictions of generally higher quality. This discrepancy arises from the differing sampling distributions of DPO data; Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of "unsure" predictions, resulting in more conservative model behavior. We will further explore the differences between these two approaches in subsequent sections.
Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority. However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as "unsure." This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score. In extreme cases, a model could assign "unsure" to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
5.3 Refinement for Unsure Prediction
We trained an automatic refinement model for each dataset as described in Section 4.3. We refine the predictions labeled as "unsure" from the initial stage and merge the refined outputs with the "sure" predictions to generate the final results. The overall evaluation results are presented in Table 2. For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Prompt Direct 78.86 36.01 82.40 36.46 32.41 13.58 Verbose 76.80 36.69 76.93 27.99 23.11 9.00 SR 78.68 29.72 76.50 30.74 29.90 11.25 Uncertainty Prob 75.89 24.19 73.88 24.86 28.48 9.63 SC 70.98 20.70 69.98 24.86 25.60 4.39 P(True) 73.88 19.18 67.07 19.03 26.29 7.34 Reliability (Ours) SFT 85.95 53.76 85.21 60.09 74.04 27.62 + DPO Joint 86.47 54.31 84.53 61.37 75.47 29.18 + DPO Post 86.55 56.33 85.87 64.95 73.96 28.13 Qwen2.5-7B Prompt Direct 78.66 36.35 83.23 37.39 42.02 16.93 Verbose 67.59 26.45 74.37 27.38 27.79 9.19 SR 68.88 19.09 67.18 19.15 34.77 8.78 Uncertainty Prob 74.18 23.56 73.97 25.64 35.91 10.81 SC 59.34 23.45 65.60 28.31 34.34 3.06 P(True) 67.52 17.05 66.25 17.99 33.66 8.07 Reliability (Ours) SFT 84.50 51.90 85.29 58.39 74.86 28.02 + DPO Joint 87.06 52.41 85.44 62.31 74.85 27.06 + DPO Post 87.57 53.19 85.86 63.61 74.95 28.51
It is important to note that, referring to Section 5.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not ideal, excessive "unsure" predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced "sure" and "unsure" classification. An ideal model should minimize unnecessary "unsure" predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
5.4 Scalability Analysis
To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities. We regard the original LLAMA3 8B model as a lower-performance model (denoted as Low) and applied supervised fine-tuning (SFT) on a subset of the dataset to obtain models with improved performance. We trained two foundation models using 1,000 samples (Medium) and 10,000 samples (High) and compared their performance against baseline algorithms, with results shown in Figure 3. For RQ3, our methods consistently outperform baselines across different foundation models after refinement. Detailed results are available in Appendix C, leading to consistent conclusions.
Additionally, as illustrated in Appendix Table 4, the advantage of our DPO Post Training method over Joint Training increases with the improvement of foundation model performance. This is due to the offline nature of our DPO algorithm; as foundational model performance improves, the limitations of the Joint Training data distribution relative to Post Training become more pronounced, impacting overall performance. Notably, in the SGD dataset, DPO Joint Training becomes less beneficial as the foundation model’s performance rises.
5.5 Cost Analysis
In this work, "Cost" refers to the additional overhead from the refinement process. We define the cost of refining an "unsure" prediction as a constant value of 1, equating Cost to the number of "unsure" predictions. For better comparison, we express Cost as the proportion of "unsure" predictions.
We compare the performance and Cost of various baselines, as shown in Figure 4. Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix D. Our methods demonstrate significant performance improvements while maintaining a relatively low Cost. Furthermore, as illustrated in Appendix Table 3, Cost decreases significantly with improved foundation model performance. This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
5.6 DPO Preference Strategies
We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in Section 3.2.1. Weighted-F1 includes two adjustable parameters: influences Weighted Precision, while affects Weighted Recall. Using the LLAMA3 8B Reliability-SFT model, we generated DPO preference data with various values and conducted Post Training. Test results on the MultiWOZ-2.4 dataset are shown in Figures 5(a) and 5(b). The model achieves optimal performance with set to 0.25 and to 0.75, reaching near-best results at a relatively low cost.
Analysis indicates that increasing makes Weighted Precision more sensitive to errors in "unsure" predictions, reducing tolerance and generating fewer "unsure" outputs. In contrast, increasing raises the reward for successful recall from "unsure" predictions, encouraging the model to produce more "unsure" predictions and adopt a more conservative stance. Practically, the optimal alpha settings depend on the scenario and refinement model performance. When the refinement model is underperforming, it may be unwise to incur significant costs for marginal performance gains. The flexible adjustment of both parameters allows for adaptation to diverse real-world situations.
5.7 Refinement Model comparison
We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 6. In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series DeepSeek-AI et al. (2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities. The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o. Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
6 Conclusion
In this work, we propose a two-stage framework that enhances the reliability of large language models by integrating fast and slow thinking paradigms. Through precise confidence assessment and high-accuracy refinement, we achieve a balance between reliability and usability. Extensive evaluations demonstrate that our method significantly improves model self-awareness and performance across complex tasks, establishing a new paradigm for enhancing the reliability of language models in error-sensitive applications.
7 Limitations
While our EKBM framework demonstrates promising results, several limitations warrant discussion. First, the refinement stage incurs computational overhead proportional to the volume of "unsure" predictions, which may hinder scalability in latency-sensitive applications. Second, the efficacy of refinement critically depends on the performance of the secondary model; suboptimal refinement can propagate errors, offsetting gains from uncertainty labeling. Third, the manual tuning of Weighted-F1 parameters (, ) introduces scenario-specific biases, limiting generalizability. Finally, our evaluation focuses on dialogue state tracking tasks—broader validation across diverse domains is necessary to assess universal applicability. Addressing these challenges could further enhance the framework’s practicality and robustness.
Appendix A Refinement Model
We trained an automated refinement model for each dataset. The refinement model conducts in-depth reasoning on unsure predictions based on task-specific criteria, operating within a defined action space. Our refinement models are built upon SFT of LLAMA3 8B utilizing data constructed with OpenAI’s GPT-4o.
Action Space
The action space of the refinement model includes three actions: remove, reserve, and correct. A prediction should be reserved if entirely correct, should be corrected if partially correct and should be removed if entirely unnecessary. For any unsure predictions, a proper action from these three options is sufficient to achieve the required refinement.
Data Construction
We adopt a multi-step reasoning Chain-of-Thought (CoT) paradigm. The specific reasoning instructions and processes are detailed in Table 5. For data construction, we used OpenAI’s GPT-4o model. During the data generation process, we provided the model with the prediction to be refined and the corresponding golden label, then instructed it to generate a reasoning process based on predefined principles, as shown in Table 6. We combined the instruction, input, the prediction to be refined, the thinking steps, and the refinement result into a final training sample. For each dataset, we sampled 5,000 examples from the model’s unsure predictions as well as from randomly selected data in the training set, ensuring a balanced distribution of data across three actions.
Appendix B Baselines Details
Here we detail the specifics of the baselines.
Prompt-based
Direct: Utilizes prompts similar to those in Table 8 to guide the model in generating predictions directly, without incorporating confidence labels.
Verbose: Employs prompts akin to those in Figure 8 to direct the model, which outputs confidence labels (sure or unsure) alongside predictions based on its self-assessment.
Self-Reflection (SR): Builds on predictions from the Direct baseline, allowing the model to reflect on each prediction to assess its confidence level.
Uncertainty-based
Token Probability (Prob): Computes average token-level probabilities, classifying outputs as "sure" or "unsure" based on a threshold , with probabilities normalized and set to 0.8.
Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as "sure," while others are marked as "unsure." We conduct ten samples () and set to 50%.
P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding "True" than "False" as "sure," otherwise as "unsure."
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA Qwen LLAMA Qwen LLAMA Qwen Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Verbose 87.28 30.62 80.69 38.05 90.30 74.07 SR 55.43 28.07 95.50 8.55 84.77 15.07 Uncertainty Prob 36.47 40.65 50.46 52.91 56.68 53.01 SC 69.26 58.91 58.05 62.59 96.80 92.71 P(True) 14.63 10.22 13.35 1.82 1.52 1.71 Reliability (Ours) SFT 32.67 30.98 20.03 21.78 51.24 55.96 + DPO Joint 31.96 31.98 21.19 23.63 57.72 45.29 + DPO Post 27.69 28.38 10.54 18.14 40.68 53.10
Appendix C Scalability Analysis
The detailed results are presented in Table 4, demonstrating that our method maintains superiority across various capability baselines, achieving more reliable outcomes (higher , , and Quality-F1) and enhanced overall performance post-refinement (higher JGA).
Method Type Method MultiWOZ-2.4 BiTOD SGD 1000Samples-Trained LLAMA3 Based Prompt Direct 91.07 86.28 87.84 42.73 93.01 92.78 92.82 70.91 82.28 85.10 82.77 38.13 Uncertainty Prob 91.84 86.52 86.84 45.43 94.93 93.55 93.66 72.73 84.45 85.24 82.80 39.48 SC 92.68 92.37 84.31 43.18 95.70 95.12 92.05 70.37 86.61 92.30 78.71 28.62 Reliability (Ours) SFT 95.42 92.11 89.86 55.62 94.97 94.74 93.80 73.64 85.16 85.14 81.06 39.23 + DPO Joint 95.11 90.68 89.05 58.94 96.06 95.60 95.11 75.36 81.66 82.82 79.69 37.62 + DPO Post 96.33 92.15 89.91 59.12 96.22 95.54 95.43 77.70 86.67 85.44 83.27 41.48 10000Samples-Trained LLAMA3 Based Prompt Direct 94.51 93.30 93.63 63.19 95.35 95.14 94.34 80.76 83.24 83.75 83.11 43.96 Uncertainty Prob 95.77 94.06 92.95 67.39 95.95 95.53 95.23 81.26 84.37 84.61 83.08 43.52 SC 96.24 96.40 92.13 67.61 97.83 97.81 93.42 78.04 88.13 92.42 79.26 38.20 Reliability (Ours) SFT 96.98 95.86 94.40 70.85 96.44 96.20 96.03 80.54 85.30 83.92 83.57 45.11 + DPO Joint 96.98 95.61 94.71 70.84 97.15 96.65 95.96 82.35 82.31 82.36 80.75 40.48 + DPO Post 97.71 96.32 94.80 71.51 96.57 96.35 96.59 82.69 86.29 85.53 84.21 45.43
| Refinement Examples |
| Give you a dialog history between USER and SYSTEM, and a slot-value pair extracted from the dialog, I want you to analyze on the dialog history and review/refine the given dialog state slot-value pair to make them more accurate and reliable. ## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Question Dialog history: USER: Hello. I need train to London liverpool Street. SYSTEM: Where are you departing from and do you have a time preference? USER: Yes, I’d like to leave Cambridge sometime after 7:00. SYSTEM: I have one that leaves at 7:59 and 4 more that depart every two hours after. USER: The 7:59 will be fine. Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: "leaveat" is one of the slots in the train domain and the value is incorrectly specified as "7:59" (formation error). Precise time should strictly follow the "HH:MM" format. step2: semantic aspect: "leaveat" is not a "name"-related slot, so the value must be extracted from the user’s direct intent. Although the system provides a train that departs at 7:59, the user only mentions that he’d like to leave Cambridge after 7:00. Considering the formation, the correct value should be 07:00. step3: domain train correct; slot leaveat correct; value 7:59 incorrect. step4: the predicted slot-value pair is incorrect in both structure and semantic, and the value should be corrected. Refine action: correct Refined slot-value pair: train: leaveat = 07:00 |
| GPT4o Prompt For Refinement Thinking Steps Generation |
| Give you a dialog history between USER and SYSTEM, a predicted slot-value pair and a gold refinement, I want you to analyze on the dialog history, refer to the examples I’ve provided and the golden refinement, and then continue to provide the thinking steps following the predefined principles. Firstly, we use json dict to describe the slots and their corresponding value space. Then, we will specify the requirements you need to comply and provide some examples. Last, we will present you with the dialog history, the predicted dialog state slot-value pair and the gold refinement for you to perform continuation. ## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Meta Requirements 1. Analyze the dialog history and the predicted slot-value pair carefully, correct and refine the pair according to the following Detailed Refinement Principles. 2. For the given slot-value pair, you should only take one of the actions of "reserve", "remove" or "correct" according to the following principles. a. reserve: if the slot is relevant to the dialog history and the value is correctly specified, you should keep the pair. b. remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print "none" indicating that the pair has been removed. c. correct: if the slot is relevant to the dialog history and the value is incorrectly specified, you should extract the correct value and replace it. 3. You need to think and reason in the same way as the Examples. Please refer to the Examples for formatting requirement. Make sure the output is all lowercase. Only focus on the given slot-value pair and do not consider any other slot. only provide the reasonable thinking steps, do not provide any extra prefixes or suffixes. ## Detailed Refinement Principles ## Examples ## Question Dialog history: USER: … SYSTEM: … Predicted slot-value pair: train: leaveat = 7:59 Refinement: Refine action: correct Refined slot-value pair: train: leaveat = 07:00 Thinking steps: |
Appendix D Refinement Cost
The detailed overall results of cost are presented in Table 3. The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 7 and Figure 8, respectively.
Appendix E LLM Instructions
The instructions used in this work are presented in this chapter, taking the MultiWOZ-2.4 dataset as an example. The instructions for the BiTOD dataset are similar.
E.1 Inference Instruction
The prompts used with the original untrained LLAMA3 model are shown in Tables 7 and 8. Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
| Direct Inference Instruction |
| Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as "dontcare" if user doesn’t have a preference. 3. You only need to consider the domains and slots that are relevant to the conversation history. Do not include those irrelevant in your response and avoid presenting empty domains or slots. 4. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. ## Output Dialog State Example {"hotel": {"area": "centre", "name": "alexander bed and breakfast", "parking": "yes", "type": "guesthouse"}, "attraction": {"name": "kambar"}} ## Examples shots… |
| Reliability Inference Instruction |
| Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as "dontcare" if user doesn’t have a preference. 3. You should try to cover as many domians and slot-value pairs relevant to the conversation history as possible. Mark each slot with either "sure" or "unsure" according to your confidence in it. For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with "sure". Otherwise, you should tag the slot-value pair with "unsure". 4. Principles: a. Goal one: Achieving as close to 100% accuracy as possible in those "sure" slot-value pairs. b. Goal two: The "sure" part along with the "unsure" part should cover all possible slots involved in the dialog history as much as possible. c. Heavy penalization: Providing incorrect slot-value pairs in the "sure" part. d. Heavy penalization: Missing slot-value pairs that should be extracted. e. Light penalization: Providing incorrect or redundant slot-value pairs in the "unsure" part. 5. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. Please refer to the Output Dialog State Example for formatting requirement. ## Output Dialog State Example {"hotel": {"area": {"value": "centre", "confidence": "sure"}, "name": {"value": "alexander bed and breakfast", "confidence": "sure"}, "parking": {"value": "yes", "confidence": "unsure"}, "type": {"value": "guesthouse", "confidence": "sure"}}, "attraction": {"name": {"value": "kambar", "confidence": "unsure"}}} ## Examples shots… |
E.2 Training Instruction
The prompts used with the trained LLAMA3 model are shown in Tables 9 and 10. The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results. In contrast, the Reliability Training Instruction is designed to train models capable of explicitly outputting confidence labels alongside predictions.
| Direct Training Instruction |
| Generate the dialogue state based on the given dialogue context. |
| Reliability Training Instruction |
| Generate the dialogue state based on the given dialogue context. Ensure the results are as reliable as possible, the ’sure’ parts are as accurate as possible, and the overall coverage includes all relevant slots. |

### v3

@@
 Abstract
-Large language models (LLMs) frequently hallucinate due to misaligned self-awareness, generating erroneous outputs when addressing queries beyond their knowledge boundaries.
-While existing approaches mitigate hallucinations via uncertainty estimation or query rejection, they suffer from computational inefficiency or sacrificed helpfulness.
+Large language models (LLMs) are prone to hallucination stemming from misaligned self-awareness, particularly when processing queries exceeding their knowledge boundaries.
+While existing mitigation strategies employ uncertainty estimation or query rejection mechanisms, they suffer from computational efficiency and sacrificed helpfulness.
 To address these issues, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework, integrating fast and slow reasoning systems to harmonize reliability and usability.
-The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate use of high-confidence outputs.
-For uncertain predictions, a slow refinement model conducts targeted reasoning to improve accuracy.
+The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate utilization of high-confidence outputs, whereas uncertain predictions trigger a slow refinement model for accuracy improvement.
 To align model behavior with our proposed object, we propose a hybrid training pipeline, enhancing self-awareness without degrading task performance.
@@
 Further analysis reveals that refinement substantially boosts accuracy while maintaining low computational overhead.
-Our work establishes a scalable paradigm for advancing LLM reliability and balancing accuracy and practical utility in error-sensitive applications.
-Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling
-Hang Zheng1††thanks: Equal contribution., Hongshen Xu1, Yuncong Liu1, Lu Chen1, Pascale Fung2, Kai Yu1 1X-LANCE Lab, Department of Computer Science and Engineering, Shanghai Jiao Tong University 2Center for Artificial Intelligence Research (CAiRE), Hong Kong University of Science and Technology {azure123, xuhongshen, chenlusz, kai.yu}@sjtu.edu.cn
+The framework establishes a scalable paradigm for deploying reliable LLMs in error-sensitive applications, effectively balancing accuracy and practical utility.
 1 Introduction
-Recently, large language models (LLMs) have exhibited impressive text generation capabilities Abdullah et al.
-(2022).
-However, despite their advancements, LLMs are prone to hallucination, where generated content misaligns with context or factual information Zhang et al.
-(2023).
-Such inaccuracies can be particularly harmful in low error-tolerance applications, eroding user trust in LLM reliability.
-Extensive works have been devoted to mitigating hallucinations in LLMs Tonmoy et al.
-(2024).
-Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, leading to errors when models attempt to answer questions beyond their expertiseLi et al.
-(2024b).
+Recently, large language models (LLMs) have demonstrated impressive text generation capabilities (Abdullah et al., 2022).
+However, LLMs are susceptible to hallucinations, where generated content misaligns with context or factual information (Zhang et al., 2023).
+Hallucinations can be particularly detrimental in applications with low tolerance for error, eroding user trust in LLM reliability.
+Extensive works have been devoted to mitigating hallucinations in LLMs (Tonmoy et al., 2024).
+Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, particularly when when addressing queries beyond its expertise (Li et al., 2024b).
 Improving the model’s self-awareness—its capacity to accurately assess its own knowledge boundaries and outputs correctness, can effectively mitigate this misalignment.
-Several uncertainty-based methods have been proposed to estimate model confidence and indirectly reflect the model’s self-awareness Huang et al.
-(2024).
-However, these approaches are not inherent to the model’s endogenous capabilities and frequently suffer from high computational costs or instability across thresholds.
-Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries Xu et al.
-(2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
-To mitigate these limitations, we propose a novel alignment objective: improving model awareness for knowledge boundaries, enabling models to respond maximally while explicitly distinguishing outputs of high and low confidence.
-We introduce the concepts of "sure" and "unsure" to clarify this distinction.
-A reliable system should deliver near-perfect accuracy for "sure" predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
+Uncertainty-based methods estimate model confidence and indirectly reflect the model’s self-awareness (Huang et al., 2024), facing challenges of inherence to the model’s endogenous capabilities, frequently sufferring from high computational costs or instability across thresholds.
+Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries (Xu et al., 2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
+To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond maximally while explicitly distinguishing high/low confidence outputs.
+We introduce the concepts of ”sure” and ”unsure” to clarify this distinction.
+A reliable system should deliver near-perfect accuracy for ”sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
 Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework.
-EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing immediate utilization of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to enhance accuracy of low-confidence predictions.
+EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions.
 Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage.
-The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective operation.
+The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective functioning.
 To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness.
-Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately categorize its outputs into "sure" and "unsure" categories compared to baseline methods.
-Additionally, we develop an automatic refinement model for "unsure" outputs, showing that the EKBM system achieve superior performance compared to traditional supervised fine-tuning (SFT) approaches.
-Our analysis of various training methods and model behaviors provides further valuable insights.
+Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately distinguish the confidence levels of its output compared to baseline methods.
+Additionally, we develop an automatic refinement model for ”unsure” outputs.
+Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
 Our contributions are summarized as follows:
@@
 We perform extensive experiments to demonstrate the effectiveness and scalability of the EKBM framework, providing valuable insights for future research.
-2 Related Work
-The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs Yin et al.
-(2024); Li et al.
-(2024b).
-Misalignment between model behavior and knowledge boundaries can result in factual hallucinations Huang et al.
-(2023) and ambiguous responses Liu et al.
-(2023).
-Various methods have been proposed to assess these boundaries.
-Uncertainty-based approaches quantify prediction confidence through token probabilities Manakul et al.
-(2023) and consistency Chen and Mueller (2024).
-Calibration strategies, including prompting Tian et al.
-(2023) and fine-tuning Tao et al.
-(2024) for improved confidence expression, align model confidence with prediction accuracy.
-Internal State Probing evaluates prediction factuality by analyzing model states, such as activations Li et al.
-(2024a).
-While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes Huang et al.
-(2024); Zhou et al.
-(2024b).
-To address this misalignment, researchers have implemented knowledge-enhanced fine-tuning Zhou et al.
-(2024a) and retrieval-augmented generation (RAG, Lewis et al., 2020) to bolster model capabilities, thereby reducing hallucinations.
-Some approaches also guide models in self-reflection to minimize self-contradictions Wei et al.
-(2022); Ji et al.
-(2023), albeit at the cost of increased overhead.
-For queries beyond the model’s capabilities, refusal or rejection strategies can prevent misinformation Xu et al.
-(2024b); Chen et al.
-(2024), though they may inadvertently reduce overall helpfulness.
 2 Problem Formulation
@@
 where denotes a scoring function, is the input prompt, and and denote correct and incorrect responses, respectively, ensuring the score of correct responses exceeds that of incorrect ones.
-Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries.
-For unanswerable queries, the model is encouraged to refrain from responding.
+Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries, where model is encouraged to directly reject the unanswerable ones.
 The corresponding alignment objective can be articulated as:
@@
 While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios.
-Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence ("sure") and low-confidence ("unsure") responses.
-The model must ensure high accuracy in the "sure" category, while tolerating some errors in the "unsure" category, which can serve as references for users or inputs for further refinement.
+Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence (”sure”) and low-confidence (”unsure”) responses.
+The model must ensure high accuracy in the ”sure” category, while tolerating some errors in the ”unsure” category, which can serve as references for users or inputs for further refinement.
 Our proposed alignment objective is formalized as:
-Here, and represent correct and incorrect responses, respectively, while and represent "sure" and "unsure" as levels of confidence.
-This objective guarantees that correct "sure" predictions are prioritized, followed by correct "unsure" predictions, which, although less confident, retain utility.
-Incorrect predictions in the "unsure" category are discouraged but more acceptable than those in the "sure" category, as the latter significantly undermine reliability.
+Here, and represent correct and incorrect responses, respectively, while and represent ”sure” and ”unsure” as levels of confidence.
+This objective guarantees that correct ”sure” predictions are prioritized, followed by correct ”unsure” predictions, which, while less confident, still provide utility.
+Incorrect predictions in the ”unsure” category are discouraged but deemed more acceptable than those in the ”sure” category, as the latter significantly undermine overall reliability.
 2.2 Reliability Evaluation
@@
 where TP, FP and FN represent True Positives, False Positives and False Negatives respectively.
-We modify the F1 score to adapt to the alignment objective of our EKBM framework.
-Since predictions labeled as "unsure" represent an intermediate state of uncertainty, we assign a weighted contribution (denoted by and ) to predictions in this category.
-The weights reflect the partial utility of "unsure" predictions as potential candidates for refinement.
+We modify the F1 score to align with the objectives of our EKBM framework.
+Since ”unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement.
 The modified precision and recall are defined as:
-Where STP and SFP represent Sure True Positives and Sure False Positives, while UTP and UFP denote the corresponding cases of Unsure.
-For UTP in Recall calculation, we only require that a prediction is partially correct (able to be refined).
+where STP and SFP denote Sure True Positives and Sure False Positives, while UTP and UFP represent the corresponding cases in the Unsure category.
+For UTP in the recall calculation, we only require that a prediction be partially correct (i.e., refinable).
+Unlike precision, we do not introduce the weighting parameter in the denominator of recall.
+This is because recall measures the proportion of correct predictions relative to the total number of gold labels, which remains unchanged regardless of the introduction of .
+Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1].
 The modified Weighted-F1 score is defined as:
-The design allows flexible handling of "unsure" predictions, effectively managing their intermediate utility.
-influences the calculation of Precision, reflecting the model’s tolerance for errors among "unsure" predictions; as increases, tolerance for such errors decreases.
-Conversely, affects Recall, indicating the degree of reward for providing additional reference information for "unsure" predictions.
-As increases, the model is encouraged to enhance helpfulness by generating more "unsure" predictions as a basis for refinement.
-We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all "unsure" predictions.
-(2) = 0, = 1: Errors in the "unsure" category are ignored, rewarding effective recall, representing the theoretical upper limit post-perfect refinement.
+2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
+The Weighted-F1 metric structurally aligns with our four-tiered alignment objective (Equation (3)) through a parametric formulation that encodes hierarchical priorities.
+Its design ensures: (1) Strict enforcement for sure predictions: Unweighted sure terms prioritize high-confidence correctness (STP) while imposing maximal penalties for sure errors (SFP); (2) Controlled utility for unsure predictions: -scaled unsure terms modulate low-confidence outputs—attenuated UFP penalties tolerate unsure errors, whereas partial UTP rewards incentivize refinable correct predictions.
+This dual-parameter mechanism enables scenario-specific adaptation: Weighted-Precision regulates error tolerance via (higher tightens unsure-error constraints) and Weighted-Recall controls reference-value incentives via (higher promotes helpfulness through unsure-correct outputs).
+The composite metric thus operationalizes the reliability-helpfulness trade-off, permitting calibration from conservative (, ) to exploratory (, ) strategies.
+We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all ”unsure” predictions.
+(2) = 0, = 1: Errors in the ”unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement.
 (3) = 1, = 0: Lacks significant interpretation.
 (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
-3.3 Reliability Metrics
-We employ two specific Weighted-F1 values as performance metrics:
-Optimal-F1 Defined as , representing the theoretical upper limit after perfect refinement of "unsure" predictions.
-Quality-F1 Defined as , it assigns "unsure" predictions half the weight of "sure" ones, encouraging the model to select "sure" when confident while minimizing penalties for "unsure" outputs.
-Quality-F1 measures the model’s self-awareness in accurately classifying both types.
-Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
+2.2.3 Reliability Metrics
+We employ two specific Weighted-F1 scores as performance metrics. Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
+Quality-F1 Defined as , it assigns ”unsure” predictions half the weight of ”sure” ones, encouraging the model to select ”sure” when confident while minimizing penalties for ”unsure” outputs.
+Quality-F1 measures the model’s self-awareness ability and reliability.
+Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of ”unsure” predictions, reflecting the overall potential of the two-stage system.
 3 Method
 3.1 Explicit Knowledge Boundary Modeling
-We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the "unsure" predictions.
-As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages.
+We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the ”unsure” predictions.
+As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful.
 Fast Prediction with Confidence Labeling
-In the first stage, the model performs the primary task while simultaneously assigning a confidence label ("sure" or "unsure") to each prediction.
-Predictions labeled as "sure" are directly accepted as final outputs, ensuring high precision.
+In the first stage, the model performs the primary task while simultaneously assigning a confidence label (”sure” or ”unsure”) to each prediction.
+Predictions labeled as ”sure” are directly accepted as final outputs, ensuring high precision.
 Slow Refinement for Unsure Prediction
-In the second stage, the "unsure" predictions are further refined.
-This refinement process can involve various strategies, such as user confirmation, automated multi-step reasoning, or other post-processing techniques.
-This additional stage incurs extra cost but increases the overall accuracy and reliability of the system.
-Our two-stage framework balances immediate usability with comprehensive coverage, making the model both reliable and helpful.
+In the second stage, ”unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
 3.2 Reliability Alignment
-To align the model’s behavior with our proposed alignment objective, we explore various training approaches to enhance its intrinsic self-awareness Xu et al.
-(2025).
-To minimize interference from task performance improvements, we introduce minimal supervisory signals related to task knowledge.
-We employ Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), with the data construction process illustrated in Figure 2 (2).
+To align the model’s behavior with our proposed alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025).
+We apply Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for alignment, with the data construction process illustrated in Figure 2(2) and two detailed examples provided in Appendix §B.
 SFT Data Construction
-We assume that task-specific knowledge boundary of the model can be approximated by repeatedly sampling the model’s outputs on the training data.
-Based on this, our SFT data construction approach is intuitive.
-We conduct multiple rounds of sampling for each training sample.
-Predictions that are consistently accurate are labeled as "sure," while those exhibiting errors or omissions receive an "unsure" label.
-Importantly, we refrain from correcting erroneous outputs to avoid introducing supervisory signals that could compromise the model’s task performance.
-The number of sampling rounds is a hyperparameter denoted as i.
-A larger i results in a more conservative model, as the proportion of "unsure" labels in the dataset increases with greater i.
+We assume that the model’s task-specific knowledge boundary can be approximated by repeatedly sampling on the training data.
+Accordingly, our SFT data construction approach is straightforward: we conduct multiple sampling for each training sample.
+Predictions consistently deemed accurate are labeled ”sure,” while those with errors or omissions are marked ”unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance.
+That is, when the model produces an incorrect prediction, we simply assign a confidence label of ”unsure” without rectifying the error.
+The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
+The number of sampling rounds, denoted as i, is a hyperparameter.
+A larger i leads to a more conservative model, as the proportion of ”unsure” labels increases with greater i in the training corpus.
 In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@
 The values in Weighted-F1 significantly influence the model’s behavior and performance.
-Through experimentation (see Section 5.6), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement.
+Through experimentation (see §4.5), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement.
 Therefore, this setting will be used by default unless specified.
-To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling, aligning with the model’s output distribution.
+To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling from the model itself, aligning with the model’s output distribution.
 3.3 Refinement Model
-We train automated refinement models for each dataset to further optimize by reasoning on unsure predictions, utilizing supervised fine-tuning (SFT) based on the LLAMA3 8B model.
-Overall, our refinement process operates at a fine-grained level, with each unsure prediction being individually refined.
-We employ a multi-step reasoning paradigm similar to the Chain of Thought Wei et al.
-(2022) methodology and use GPT-4o Hurst et al.
-(2024) to generate the reasoning process.
-This process is illustrated in Figure 2 (3), and detailed information can be found in Appendix A.
+We train automated refinement models for each dataset to further optimize unsure predictions by multi-step reasoning, leveraging supervised fine-tuning (SFT) based on the LLAMA3 8B model.
+Our refinement operates at a fine-grained level, addressing each unsure prediction individually.
+We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction.
+This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §C.
+4 Experiments
+We conduct experiments focusing on three core research questions:
+RQ1 Does our reliability training pipeline outperform current methods in enhancing self-awareness?
+RQ2 Does our EKBM framework improve the overall performance of LLMs on complex tasks?
+RQ3 Does our approach scale robustly across foundation models with varying capabilities?
+4.1 Experiments Setup
+4.1.1 Model and Baselines
+We utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone for our experiments.
+For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation.
+The key distinction lies in the different source distributions of preference data sampled for DPO.
+Training details could be found in Appendix §A.
+To evaluate the reliability of our proposed models, we compare them against several baselines, including three prompt-based and three uncertainty-based approaches (see Appendix §D).
+Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models for comparison.
+By default, we utilize only 1,000 samples for Reliability-SFT training and 2,000 for DPO training—significantly less than traditional models.
+Prompt-based
+methods encompass three approaches: Direct, Verbose, and Self-Reflection (SR).
+The Direct approach generates predictions without confidence estimation, whereas Verbose simultaneously produces confidence labels.
+SR assesses each prediction through self-evaluative reflection.
+Uncertainty-based
+methods include Token Probability (Prob), Self-Consistency (SC), and P(True) (Kadavath et al., 2022), each employing distinct techniques to classify outputs as ”sure” or ”unsure”.
+Prob calculates average token probabilities, SC measures the frequency of predictions across multiple samples, and P(True) determines the probability of the ”True” token during model self-evaluation.
 Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Dubey et al.
@@
 (2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
-5 Experiments
-We conduct experiments focusing on three core research questions:
-RQ1 Does our proposed reliability training pipeline outperform current methods in enhancing self-awareness abilities?
-RQ2 Does our EKBM framework improve the overall performance of LLMs on complex tasks?
-RQ3 Does our approach demonstrate strong scalability, performing well across foundation models with various performance?
-5.1 Experiments Setup
-5.1.1 Model and Baselines
-For our experiments, we utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone.
-For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation.
-The key distinction lies in the different source distributions of preference data sampled for DPO.
-To evaluate the reliability of our proposed methods, we compare their performance against several baseline approaches, including three prompt-based and three uncertainty-based methods.
-Prompt-based
-methods include three approaches: Direct, Verbose, and Self-Reflection (SR).
-The Direct method predicts outputs without incorporating any confidence labels, while Verbose generates confidence labels alongside predictions based on predefined principles.
-SR evaluates prediction reliability through individual reflections.
-Uncertainty-based
-methods consist of three approaches: Token Probability (Prob), Self-Consistency (SC), and P(True) Kadavath et al.
-(2022).
-These methods classify outputs as "sure" or "unsure" using different techniques.
-Prob calculates average token-level probabilities, SC assesses the frequency of predictions across multiple samples, and P(True) determines the probability of the "True" token during model self-evaluation.
-The details of the baselines can be found in E.1.
 4.1.2 Dataset
-We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology Xu et al.
-(2024a).
+We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology (Xu et al., 2024a).
+We focus on the intricate multi-slot problem, where the model is likely to exhibit varying confidence levels for different parts of responses to the same query.
+Consequently, we do not adopt traditional QA or mathematical datasets.
+To ensure a robust evaluation, we utilize three datasets: MultiWOZ-2.4 (Ye et al., 2021), BiTOD (Lin et al., 2021) and SGD (Rastogi et al., 2020).
+For BiTOD, we adopt the English version, while for SGD, we randomly select 10,000 instances from the testset to mitigate excessive computational overhead.
 Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1.
 For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.
-To ensure a robust evaluation, we select three datasets: MultiWOZ-2.4 Ye et al.
-(2021), BiTOD Lin et al.
-, and SGD Rastogi et al.
-(2020).
-For BiTOD we only utilize the English version.
-For SGD, we randomly select 10,000 test samples.
-5.1.3 Training Details
-We modify the DeepSpeed-Chat Yao et al.
-(2023) framework to accommodate our experimental setup.
-We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others.
-By default, we utilize 1,000samples for Reliability-SFT Training and 2,000samples for DPO Training.
-All models are trained for 1 epoch using NVIDIA A800 GPUs.
 4.2 Reliability Evaluation
-The results of our experiments on model reliability are summarized in Table 1.
-We compare our reliability-training pipeline with baseline algorithms, emphasizing the model’s self-awareness ability.
-For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, markedly enhancing self-awareness.
-Our approach consistently achieves higher Quality-F1 scores, indicating improved effectiveness in categorizing predictions as "sure" or "unsure" based on confidence.
-Notably, a high Precisionsure indicates reliable "sure" predictions, fostering user trust with minimal verification.
-Simultaneously, a high Recalltotal ensures that the "unsure" category encompasses a broad range of potentially correct outputs, forming a solid basis for further refinement.
+The experimental results on model reliability are summarized in Table 1.
+We evaluate our reliability-training pipeline against baseline algorithms, with a particular focus on the model’s self-awareness capabilities.
+For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, demonstrating a substantial improvement in self-awareness.
+Our approach consistently achieves superior Quality-F1 scores, indicating improved effectiveness in distinguishing predictions as ”sure” or ”unsure” based on confidence levels.
+Notably, a high indicates reliable ”sure” predictions, fostering user trust with minimal need for verification.
+Simultaneously, a high ensures that the ”unsure” category encompasses a broad spectrum of potentially correct outputs, forming a solid basis for further refinement.
 Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance.
-In other words, our system offers improved performance by effectively balancing precision and recall, ensuring both immediate reliability and the potential for improved post-refinement accuracy.
+In other words, our system optimally balances precision and recall, delivering both immediate reliability and greater potential for post-refinement accuracy.
 In our approach, methods incorporating DPO generally outperform SFT-only methods.
 Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores.
-This suggests that Joint Training models are more conservative, relying on subsequent refinement, whereas Post Training models exhibit greater confidence, yielding predictions of generally higher quality.
-This discrepancy arises from the differing sampling distributions of DPO data; Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of "unsure" predictions, resulting in more conservative model behavior.
-We will further explore the differences between these two approaches in subsequent sections.
+This suggests that Joint Training models adopt a more conservative stance, deferring to subsequent refinement, whereas Post Training models exhibit greater confidence, producing predictions of generally higher quality.
+This discrepancy stems from the differing sampling distributions of DPO data: Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of ”unsure” predictions and resulting in more cautious model behavior.
 Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority.
-However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as "unsure." This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score.
-In extreme cases, a model could assign "unsure" to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
+However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as ”unsure.” This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score.
+In extreme cases, a model could assign ”unsure” to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
 4.3 Refinement for Unsure Prediction
 We trained an automatic refinement model for each dataset as described in Section 3.3.
-We refine the predictions labeled as "unsure" from the initial stage and merge the refined outputs with the "sure" predictions to generate the final results.
+We refine the predictions labeled as ”unsure” from the initial stage and merge the refined outputs with the ”sure” predictions to generate the final results.
 The overall evaluation results are presented in Table 2.
 For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
-Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Prompt Direct 78.86 36.01 82.40 36.46 32.41 13.58 Verbose 76.80 36.69 76.93 27.99 23.11 9.00 SR 78.68 29.72 76.50 30.74 29.90 11.25 Uncertainty Prob 75.89 24.19 73.88 24.86 28.48 9.63 SC 70.98 20.70 69.98 24.86 25.60 4.39 P(True) 73.88 19.18 67.07 19.03 26.29 7.34 Reliability (Ours) SFT 85.95 53.76 85.21 60.09 74.04 27.62 + DPO Joint 86.47 54.31 84.53 61.37 75.47 29.18 + DPO Post 86.55 56.33 85.87 64.95 73.96 28.13 Qwen2.5-7B Prompt Direct 78.66 36.35 83.23 37.39 42.02 16.93 Verbose 67.59 26.45 74.37 27.38 27.79 9.19 SR 68.88 19.09 67.18 19.15 34.77 8.78 Uncertainty Prob 74.18 23.56 73.97 25.64 35.91 10.81 SC 59.34 23.45 65.60 28.31 34.34 3.06 P(True) 67.52 17.05 66.25 17.99 33.66 8.07 Reliability (Ours) SFT 84.50 51.90 85.29 58.39 74.86 28.02 + DPO Joint 87.06 52.41 85.44 62.31 74.85 27.06 + DPO Post 87.57 53.19 85.86 63.61 74.95 28.51
-It is important to note that, referring to Section 5.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models.
+Method Type | Method | MultiWOZ-2.4 | BiTOD | SGD | |||
+LLAMA3-8B | |||||||
+Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 |
+Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |
+SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |
+Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 |
+SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |
+P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |
+Reliability (Ours) | SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 |
++ DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |
++ DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |
+Qwen2.5-7B | |||||||
+Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 |
+Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |
+SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |
+Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 |
+SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |
+P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |
+Reliability (Ours) | SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 |
++ DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |
++ DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 |
+It is important to note that, referring to §4.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models.
 This discrepancy stems from the limitations of the refinement model.
-Since the refinement model is not ideal, excessive "unsure" predictions can lead to residual errors after refinement, incurring significant costs without proportional gains.
-This highlights the importance of a balanced "sure" and "unsure" classification.
-An ideal model should minimize unnecessary "unsure" predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
+Since the refinement model is not ideal, excessive ”unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains.
+This highlights the importance of a balanced ”sure” and ”unsure” classification.
+An ideal model should minimize unnecessary ”unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
 4.4 Scalability Analysis
 To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities.
-We regard the original LLAMA3 8B model as a lower-performance model (denoted as Low) and applied supervised fine-tuning (SFT) on a subset of the dataset to obtain models with improved performance.
-We trained two foundation models using 1,000 samples (Medium) and 10,000 samples (High) and compared their performance against baseline algorithms, with results shown in Figure 3.
-For RQ3, our methods consistently outperform baselines across different foundation models after refinement.
-Detailed results are available in Appendix C, leading to consistent conclusions.
-Additionally, as illustrated in Appendix Table 4, the advantage of our DPO Post Training method over Joint Training increases with the improvement of foundation model performance.
-This is due to the offline nature of our DPO algorithm; as foundational model performance improves, the limitations of the Joint Training data distribution relative to Post Training become more pronounced, impacting overall performance.
-Notably, in the SGD dataset, DPO Joint Training becomes less beneficial as the foundation model’s performance rises.
-5.5 Cost Analysis
-In this work, "Cost" refers to the additional overhead from the refinement process.
-We define the cost of refining an "unsure" prediction as a constant value of 1, equating Cost to the number of "unsure" predictions.
-For better comparison, we express Cost as the proportion of "unsure" predictions.
+We consider the original LLAMA3 8B model a lower-performance baseline (denoted as Low) and applied supervised fine-tuning (SFT) on each dataset to enhance its DST capabilities.
+We train two foundation models using 1,000 samples (Medium) and 10,000 samples (High), then continually conduct reliability training and evaluate the performance of the obtained reliable LLM against baseline algorithms, with results shown in Figure 3.
+For RQ3, our methods consistently outperform baselines across different foundation models after refinement, demonstrating strong scalability—particularly in real-world scenarios where models are often domain-specific tuned and possess high task capabilities.
+Detailed results in Appendix §E further support these findings.
+Additionally, as illustrated in Appendix Table 5, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves.
+This trend stems from the offline nature of our DPO algorithm—higher-performing foundation models exacerbate the limitations of the Joint Training data distribution compared to Post Training, thereby affecting overall performance.
+Notably, in the SGD dataset, the benefits of DPO Joint Training diminish as the foundation model’s performance increases, even underperforming the SFT-only approach.
+4.5 DPO Preference Strategies
+We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1.
+Weighted-F1 incorporates two adjustable parameters: influences Weighted Precision, while affects Weighted Recall.
+Using the Reliability-SFT trained LLAMA3 8B model as backbone, we further generated DPO preference data with varying values and conducted consecutive Post Training.
+Test results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance with and , reaching near-best results at a relatively low cost.
+Analysis indicates that increasing heightens sensitivity to errors in ”unsure” predictions, reducing tolerance and yielding fewer ”unsure” outputs.
+Conversely, increasing amplifies rewards for successfully recalling ”unsure” predictions, encouraging the model to produce more ”unsure” predictions and adopt a more conservative stance.
+Practically, the optimal alpha settings depend on the scenario and refinement model performance.
+If the refinement model underperforms, excessive costs for marginal gains may not be justified.
+The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
+4.6 Detailed Analysis
+4.6.1 Prediction Types Analysis
+We further analyzed the model’s behavior after Reliability Training to evaluate its ability to make reliable confidence judgments in its outputs.
+Specifically, we examined the distribution of output categories and compared them to baseline models.
+These categories include five cases: correct and incorrect predictions classified as ”sure” or ”unsure”, along with missing predictions.
+The results are presented in Figure 5.
+Our findings highlight two key improvements: (1) Enhanced reliability—The proportion of incorrect predictions within the ”sure” category has been notably reduced, with most errors now concentrated in the ”unsure” category.
+This indicates a stronger self-assessment capability, allowing the model to better identify potentially erroneous predictions compared to baselines.
+Additionally, across all scenarios, the ”unsure” category contains fewer correct predictions than incorrect ones, demonstrating the model’s effectiveness in placing accurate outputs in the ”sure” category, a distinction baseline models struggle to achieve.
+(2) Improved recall—Our approach significantly reduces missing predictions, capturing previously omitted relevant information within the ”unsure” category.
+While some may still contain errors, this shift enhances overall system helpfulness and provides a stronger foundation for further refinement.
+4.6.2 Cost Analysis
+In this work, ”Cost” refers to the additional overhead from the refinement process.
+We define the cost of refining an ”unsure” prediction as a constant value of 1, equating Cost to the number of ”unsure” predictions.
+For better comparison, we express Cost as the proportion of ”unsure” predictions.
 We compare the performance and Cost of various baselines, as shown in Figure 6.
-Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix D.
+Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix §F.
 Our methods demonstrate significant performance improvements while maintaining a relatively low Cost.
@@
 This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
-5.6 DPO Preference Strategies
-We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in Section 3.2.1.
-Weighted-F1 includes two adjustable parameters: influences Weighted Precision, while affects Weighted Recall.
-Using the LLAMA3 8B Reliability-SFT model, we generated DPO preference data with various values and conducted Post Training.
-Test results on the MultiWOZ-2.4 dataset are shown in Figures 5(a) and 5(b).
-The model achieves optimal performance with set to 0.25 and to 0.75, reaching near-best results at a relatively low cost.
-Analysis indicates that increasing makes Weighted Precision more sensitive to errors in "unsure" predictions, reducing tolerance and generating fewer "unsure" outputs.
-In contrast, increasing raises the reward for successful recall from "unsure" predictions, encouraging the model to produce more "unsure" predictions and adopt a more conservative stance.
-Practically, the optimal alpha settings depend on the scenario and refinement model performance.
-When the refinement model is underperforming, it may be unwise to incur significant costs for marginal performance gains.
-The flexible adjustment of both parameters allows for adaptation to diverse real-world situations.
-5.7 Refinement Model comparison
+4.6.3 Refinement Model comparison
 We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 7.
-In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series DeepSeek-AI et al.
-(2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities.
+In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series (DeepSeek-AI et al., 2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities.
 The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o.
 Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
+5 Related Work
+The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs (Yin et al., 2024; Li et al., 2024b).
+Misalignment between model behavior and knowledge boundaries can result in factual hallucinations (Huang et al., 2023) and ambiguous responses (Liu et al., 2023).
+Various methods have been proposed to assess these boundaries.
+Uncertainty-based approaches quantify prediction confidence through token probabilities (Manakul et al., 2023) and consistency (Chen & Mueller, 2024).
+Calibration strategies, including prompting (Tian et al., 2023) and fine-tuning (Tao et al., 2024) for improved confidence expression, align model confidence with prediction accuracy.
+Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a).
+While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024).
+To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023), albeit with increased computational overhead.
+More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024).
+While effective in mitigating misinformation, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
 6 Conclusion
@@
 Extensive evaluations demonstrate that our method significantly improves model self-awareness and performance across complex tasks, establishing a new paradigm for enhancing the reliability of language models in error-sensitive applications.
-7 Limitations
-While our EKBM framework demonstrates promising results, several limitations warrant discussion.
-First, the refinement stage incurs computational overhead proportional to the volume of "unsure" predictions, which may hinder scalability in latency-sensitive applications.
-Second, the efficacy of refinement critically depends on the performance of the secondary model; suboptimal refinement can propagate errors, offsetting gains from uncertainty labeling.
-Third, the manual tuning of Weighted-F1 parameters (, ) introduces scenario-specific biases, limiting generalizability.
-Finally, our evaluation focuses on dialogue state tracking tasks—broader validation across diverse domains is necessary to assess universal applicability.
-Addressing these challenges could further enhance the framework’s practicality and robustness.
-Appendix A Refinement Model
+Appendix A Training Details
+We modify the DeepSpeed-Chat (Yao et al., 2023) framework to accommodate our experimental setup.
+We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others.
+All models are trained for 1 epoch using NVIDIA A800 GPUs.
+Appendix B Training Data Construction
+We provide detailed examples to illustrate our training data construction process, including model inputs and the expected outputs the model is intended to learn.
+The data construction for SFT is outlined in Table 3, while that for DPO is detailed in Table 4.
+Input |
+Dialog history: USER: Is there a modern European restaurant in the north?
+SYSTEM: I found city stop restaurant.
+They are located at Cambridge City Football Club Milton Road Chesterton.
+USER: That sounds lovely!
+SYSTEM: Would you like me to help you with reserving a table?
+USER: Maybe.
+Is that restaurant in the moderate price range?
+SYSTEM: No, it is in the expensive price range.
+Is this okay, or would you like to try another area?
+USER: How about a restaurant that serves Indian food?
+SYSTEM: There are several restaurants what part of town would you like it?
+USER: A restaurant in the north, please.
+Reliable dialog state: |
+Gold Labels |
+{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
+LLM Original Predictions |
+{’restaurant’: {’area’: ’north’, ’food’: ’european’, ’name’: ’Cambridge City Football Club Milton Road Chesterton’}} |
+Constructed SFT Sample Output |
+{’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’European’, ’confidence’: ’unsure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}, ’name’: {’value’: ’Cambridge City Football Club Milton Road Chesterton’, ’confidence’: ’unsure’}}} |
+Explanation |
+In the LLM’s original predictions, the value for area is correct, food is incorrect, name is extraneous, and pricerange is missing.
+Consequently, in the constructed SFT sample output, area is labeled as ”sure,” while the other three are marked as ”unsure,” with erroneous information left uncorrected.
+|
+Input |
+Dialog history: USER: Is there a modern European restaurant in the north?
+SYSTEM: I found city stop restaurant.
+They are located at Cambridge City Football Club Milton Road Chesterton.
+USER: That sounds lovely!
+SYSTEM: Would you like me to help you with reserving a table?
+USER: Maybe.
+Is that restaurant in the moderate price range?
+SYSTEM: No, it is in the expensive price range.
+Is this okay, or would you like to try another area?
+USER: How about a restaurant that serves Indian food?
+SYSTEM: There are several restaurants what part of town would you like it?
+USER: A restaurant in the north, please.
+Reliable dialog state: |
+Gold Labels |
+{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
+Multi-Sampling With the Foundation LLM |
+Constructed DPO Sample Output Pairs |
+Chosen {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}}} Rejected {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}}} |
+Explanation |
+Compared to the rejected sample, the chosen sample exhibits a higher recall for a correct pricerange within the ”unsure” category, aligning more closely with our alignment objectives.
+This is reflected in a higher Weighted-F1 score, justifying its selection over the rejected sample.
+|
+Appendix C Refinement Model
 We trained an automated refinement model for each dataset.
-The refinement model conducts in-depth reasoning on unsure predictions based on task-specific criteria, operating within a defined action space.
-Our refinement models are built upon SFT of LLAMA3 8B utilizing data constructed with OpenAI’s GPT-4o.
+The refinement model conducts in-depth step-by-step reasoning on unsure predictions based on task-specific criteria, operating within a defined action space.
+Our refinement models are built upon LLAMA3 8B utilizing SFT training with data constructed using OpenAI’s GPT-4o.
 Action Space
@@
 For each dataset, we sampled 5,000 examples from the model’s unsure predictions as well as from randomly selected data in the training set, ensuring a balanced distribution of data across three actions.
-Appendix B Baselines Details
+Appendix D Baselines Details
 Here we detail the specifics of the baselines.
@@
 Uncertainty-based
-Token Probability (Prob): Computes average token-level probabilities, classifying outputs as "sure" or "unsure" based on a threshold , with probabilities normalized and set to 0.8.
-Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as "sure," while others are marked as "unsure." We conduct ten samples () and set to 50%.
-P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding "True" than "False" as "sure," otherwise as "unsure."
-Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA Qwen LLAMA Qwen LLAMA Qwen Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Verbose 87.28 30.62 80.69 38.05 90.30 74.07 SR 55.43 28.07 95.50 8.55 84.77 15.07 Uncertainty Prob 36.47 40.65 50.46 52.91 56.68 53.01 SC 69.26 58.91 58.05 62.59 96.80 92.71 P(True) 14.63 10.22 13.35 1.82 1.52 1.71 Reliability (Ours) SFT 32.67 30.98 20.03 21.78 51.24 55.96 + DPO Joint 31.96 31.98 21.19 23.63 57.72 45.29 + DPO Post 27.69 28.38 10.54 18.14 40.68 53.10
-Appendix C Scalability Analysis
+Token Probability (Prob): Computes average token-level probabilities, classifying outputs as ”sure” or ”unsure” based on a threshold , with probabilities normalized and set to 0.8.
+Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as ”sure,” while others are marked as ”unsure.” We conduct ten samples () and set to 50%.
+P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding ”True” than ”False” as ”sure,” otherwise as ”unsure.”
+Appendix E Scalability Analysis
 The detailed results are presented in Table 5, demonstrating that our method maintains superiority across various capability baselines, achieving more reliable outcomes (higher , , and Quality-F1) and enhanced overall performance post-refinement (higher JGA).
 Method Type Method MultiWOZ-2.4 BiTOD SGD 1000Samples-Trained LLAMA3 Based Prompt Direct 91.07 86.28 87.84 42.73 93.01 92.78 92.82 70.91 82.28 85.10 82.77 38.13 Uncertainty Prob 91.84 86.52 86.84 45.43 94.93 93.55 93.66 72.73 84.45 85.24 82.80 39.48 SC 92.68 92.37 84.31 43.18 95.70 95.12 92.05 70.37 86.61 92.30 78.71 28.62 Reliability (Ours) SFT 95.42 92.11 89.86 55.62 94.97 94.74 93.80 73.64 85.16 85.14 81.06 39.23 + DPO Joint 95.11 90.68 89.05 58.94 96.06 95.60 95.11 75.36 81.66 82.82 79.69 37.62 + DPO Post 96.33 92.15 89.91 59.12 96.22 95.54 95.43 77.70 86.67 85.44 83.27 41.48 10000Samples-Trained LLAMA3 Based Prompt Direct 94.51 93.30 93.63 63.19 95.35 95.14 94.34 80.76 83.24 83.75 83.11 43.96 Uncertainty Prob 95.77 94.06 92.95 67.39 95.95 95.53 95.23 81.26 84.37 84.61 83.08 43.52 SC 96.24 96.40 92.13 67.61 97.83 97.81 93.42 78.04 88.13 92.42 79.26 38.20 Reliability (Ours) SFT 96.98 95.86 94.40 70.85 96.44 96.20 96.03 80.54 85.30 83.92 83.57 45.11 + DPO Joint 96.98 95.61 94.71 70.84 97.15 96.65 95.96 82.35 82.31 82.36 80.75 40.48 + DPO Post 97.71 96.32 94.80 71.51 96.57 96.35 96.59 82.69 86.29 85.53 84.21 45.43
+Appendix F Refinement Cost
+The detailed overall results of cost are presented in Table 7 and 7.
+The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 8(a) and Figure 8(b), respectively.
+Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA Qwen LLAMA Qwen LLAMA Qwen Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Verbose 87.28 30.62 80.69 38.05 90.30 74.07 SR 55.43 28.07 95.50 8.55 84.77 15.07 Uncertainty Prob 36.47 40.65 50.46 52.91 56.68 53.01 SC 69.26 58.91 58.05 62.59 96.80 92.71 P(True) 14.63 10.22 13.35 1.82 1.52 1.71 Reliability (Ours) SFT 32.67 30.98 20.03 21.78 51.24 55.96 + DPO Joint 31.96 31.98 21.19 23.63 57.72 45.29 + DPO Post 27.69 28.38 10.54 18.14 40.68 53.10
+Method Type Method MultiWOZ-2.4 BiTOD SGD Medium High Medium High Medium High Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Uncertainty Prob 6.55 5.74 3.29 1.37 8.14 5.53 SC 28.93 12.17 6.25 2.12 41.68 41.42 Reliability (Ours) SFT 13.52 5.12 2.16 0.66 12.60 1.87 + DPO Joint 10.70 3.94 1.36 1.04 11.66 2.67 + DPO Post 13.32 8.59 3.44 1.46 19.01 4.19
 Refinement Examples |
 Give you a dialog history between USER and SYSTEM, and a slot-value pair extracted from the dialog, I want you to analyze on the dialog history and review/refine the given dialog state slot-value pair to make them more accurate and reliable.
-## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Question Dialog history: USER: Hello.
+## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Question Dialog history: USER: Hello.
 I need train to London liverpool Street.
@@
 USER: The 7:59 will be fine.
-Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: "leaveat" is one of the slots in the train domain and the value is incorrectly specified as "7:59" (formation error).
-Precise time should strictly follow the "HH:MM" format.
-step2: semantic aspect: "leaveat" is not a "name"-related slot, so the value must be extracted from the user’s direct intent.
+Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: ”leaveat” is one of the slots in the train domain and the value is incorrectly specified as ”7:59” (formation error).
+Precise time should strictly follow the ”HH:MM” format.
+step2: semantic aspect: ”leaveat” is not a ”name”-related slot, so the value must be extracted from the user’s direct intent.
 Although the system provides a train that departs at 7:59, the user only mentions that he’d like to leave Cambridge after 7:00.
@@
 Last, we will present you with the dialog history, the predicted dialog state slot-value pair and the gold refinement for you to perform continuation.
-## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Meta Requirements 1.
+## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Meta Requirements 1.
 Analyze the dialog history and the predicted slot-value pair carefully, correct and refine the pair according to the following Detailed Refinement Principles.
 2.
-For the given slot-value pair, you should only take one of the actions of "reserve", "remove" or "correct" according to the following principles.
+For the given slot-value pair, you should only take one of the actions of ”reserve”, ”remove” or ”correct” according to the following principles.
 a.
@@
 b.
-remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print "none" indicating that the pair has been removed.
+remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print ”none” indicating that the pair has been removed.
 c.
@@
 ## Detailed Refinement Principles ## Examples ## Question Dialog history: USER: … SYSTEM: … Predicted slot-value pair: train: leaveat = 7:59 Refinement: Refine action: correct Refined slot-value pair: train: leaveat = 07:00 Thinking steps: |
-Appendix D Refinement Cost
-The detailed overall results of cost are presented in Table 3. The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 7 and Figure 8, respectively.
-Appendix E LLM Instructions
+Appendix G LLM Instructions
 The instructions used in this work are presented in this chapter, taking the MultiWOZ-2.4 dataset as an example. The instructions for the BiTOD dataset are similar.
-E.1 Inference Instruction
-The prompts used with the original untrained LLAMA3 model are shown in Tables 7 and 8.
+G.1 Inference Instruction
+The prompts used with the original untrained LLAMA3 model are shown in Tables 10 and 11.
 Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
@@
 Last, we demonstrate some use cases.
-## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1.
+## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1.
 Analyze the dialog history carefully and fill the relevant domains and slots.
@@
 For slots without specified value range, the answer must be extracted from the history.
-Set value as "dontcare" if user doesn’t have a preference.
+Set value as ”dontcare” if user doesn’t have a preference.
 3.
@@
 Do not provide any extra prefixes or suffixes or any explanations.
-## Output Dialog State Example {"hotel": {"area": "centre", "name": "alexander bed and breakfast", "parking": "yes", "type": "guesthouse"}, "attraction": {"name": "kambar"}} ## Examples shots… |
+## Output Dialog State Example {”hotel”: {”area”: ”centre”, ”name”: ”alexander bed and breakfast”, ”parking”: ”yes”, ”type”: ”guesthouse”}, ”attraction”: {”name”: ”kambar”}} ## Examples shots… |
 Reliability Inference Instruction |
@@
 Last, we demonstrate some use cases.
-## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1.
+## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1.
 Analyze the dialog history carefully and fill the relevant domains and slots.
@@
 For slots without specified value range, the answer must be extracted from the history.
-Set value as "dontcare" if user doesn’t have a preference.
+Set value as ”dontcare” if user doesn’t have a preference.
 3.
 You should try to cover as many domians and slot-value pairs relevant to the conversation history as possible.
-Mark each slot with either "sure" or "unsure" according to your confidence in it.
-For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with "sure".
-Otherwise, you should tag the slot-value pair with "unsure".
+Mark each slot with either ”sure” or ”unsure” according to your confidence in it.
+For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with ”sure”.
+Otherwise, you should tag the slot-value pair with ”unsure”.
 4.
 Principles: a.
-Goal one: Achieving as close to 100% accuracy as possible in those "sure" slot-value pairs.
+Goal one: Achieving as close to 100% accuracy as possible in those ”sure” slot-value pairs.
 b.
-Goal two: The "sure" part along with the "unsure" part should cover all possible slots involved in the dialog history as much as possible.
+Goal two: The ”sure” part along with the ”unsure” part should cover all possible slots involved in the dialog history as much as possible.
 c.
-Heavy penalization: Providing incorrect slot-value pairs in the "sure" part.
+Heavy penalization: Providing incorrect slot-value pairs in the ”sure” part.
 d.
@@
 e.
-Light penalization: Providing incorrect or redundant slot-value pairs in the "unsure" part.
+Light penalization: Providing incorrect or redundant slot-value pairs in the ”unsure” part.
 5.
@@
 Please refer to the Output Dialog State Example for formatting requirement.
-## Output Dialog State Example {"hotel": {"area": {"value": "centre", "confidence": "sure"}, "name": {"value": "alexander bed and breakfast", "confidence": "sure"}, "parking": {"value": "yes", "confidence": "unsure"}, "type": {"value": "guesthouse", "confidence": "sure"}}, "attraction": {"name": {"value": "kambar", "confidence": "unsure"}}} ## Examples shots… |
-E.2 Training Instruction
-The prompts used with the trained LLAMA3 model are shown in Tables 9 and 10.
+## Output Dialog State Example {”hotel”: {”area”: {”value”: ”centre”, ”confidence”: ”sure”}, ”name”: {”value”: ”alexander bed and breakfast”, ”confidence”: ”sure”}, ”parking”: {”value”: ”yes”, ”confidence”: ”unsure”}, ”type”: {”value”: ”guesthouse”, ”confidence”: ”sure”}}, ”attraction”: {”name”: {”value”: ”kambar”, ”confidence”: ”unsure”}}} ## Examples shots… |
+G.2 Training Instruction
+The prompts used with the trained LLAMA3 model are shown in Tables 12 and 13.
 The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results.

### v4

@@
 Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries (Xu et al., 2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
-To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond maximally while explicitly distinguishing high/low confidence outputs.
-We introduce the concepts of ”sure” and ”unsure” to clarify this distinction.
-A reliable system should deliver near-perfect accuracy for ”sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
+To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond more effectively by explicitly distinguishing between high-confidence (“sure”) and low-confidence (“unsure”) outputs.
+A reliable system should deliver near-perfect accuracy for “sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
 Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework.
 EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions.
-Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage.
-The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective functioning.
+Crucially, EKBM does not merely filter uncertain outputs but
+leverages them as opportunities for improvement, ensuring both precision and recall.
+The collaboration between these two systems strikes a balance between reliability and efficiency, with the fast-thinking model’s self-awareness being crucial for the system’s effective functioning.
 To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness.
-Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately distinguish the confidence levels of its output compared to baseline methods.
-Additionally, we develop an automatic refinement model for ”unsure” outputs.
+Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively and reliably distinguishes the confidence levels of model output compared to baseline methods.
+Additionally, we develop an automatic refinement model for “unsure” outputs.
 Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
 Our contributions are summarized as follows:
-We present the EKBM framework, integrating fast-thinking confidence labeling with slow-thinking refinement, along with metrics for evaluating LLM self-awareness and reliability.
+We present the EKBM framework, integrating fast-thinking assessment with slow-thinking refinement, along with metrics for evaluating self-awareness and reliability.
 We develop a comprehensive training pipeline that enhances LLM self-awareness, yielding more reliable models.
@@
 To enhance the reliability of LLMs, a prevailing method involves aligning their performance with high-quality training data. This alignment aims to achieve the following objective:
-where denotes a scoring function, is the input prompt, and and denote correct and incorrect responses, respectively, ensuring the score of correct responses exceeds that of incorrect ones.
+where a scoring function ensures that the score of a correct response to an input prompt is higher than that of an incorrect response .
 Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries, where model is encouraged to directly reject the unanswerable ones.
@@
 where denotes a truthful rejection, and an incorrect response. This suggests that refusal is preferable to providing an incorrect answer.
-While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios.
-Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence (”sure”) and low-confidence (”unsure”) responses.
-The model must ensure high accuracy in the ”sure” category, while tolerating some errors in the ”unsure” category, which can serve as references for users or inputs for further refinement.
+While this rejection strategy ensures reliability, it may unduly limit the model’s helpfulness.
+Therefore, we propose a more nuanced objective: the model should provide an answer whenever possible, categorizing its output as either high-confidence (“sure”) or low-confidence (“unsure”).
+The model must ensure high accuracy for “sure” responses while tolerating some errors in the “unsure” category, which can serve as references for users or inputs for further refinement.
 Our proposed alignment objective is formalized as:
-Here, and represent correct and incorrect responses, respectively, while and represent ”sure” and ”unsure” as levels of confidence.
-This objective guarantees that correct ”sure” predictions are prioritized, followed by correct ”unsure” predictions, which, while less confident, still provide utility.
-Incorrect predictions in the ”unsure” category are discouraged but deemed more acceptable than those in the ”sure” category, as the latter significantly undermine overall reliability.
+Here, and denote correct and incorrect responses, while and represent “sure” and “unsure” confidence levels.
+This objective guarantees that correct “sure” predictions are prioritized over correct “unsure” predictions, which, while less confident, still provide utility.
+Incorrect “unsure” predictions are discouraged but considered more acceptable than incorrect “sure” predictions, as the latter significantly undermine overall reliability.
 2.2 Reliability Evaluation
@@
 2.2.1 Weighted-F1
-F1-score is usually taken as the evaluation metric for multi-slot problems:
-where TP, FP and FN represent True Positives, False Positives and False Negatives respectively.
+The F1-score, calculated using True Positives (TP), False Positives (FP), and False Negatives (FN), is typically the evaluation metric for multi-slot problems:
 We modify the F1 score to align with the objectives of our EKBM framework.
-Since ”unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement.
+Since “unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement.
 The modified precision and recall are defined as:
-where STP and SFP denote Sure True Positives and Sure False Positives, while UTP and UFP represent the corresponding cases in the Unsure category.
-For UTP in the recall calculation, we only require that a prediction be partially correct (i.e., refinable).
-Unlike precision, we do not introduce the weighting parameter in the denominator of recall.
-This is because recall measures the proportion of correct predictions relative to the total number of gold labels, which remains unchanged regardless of the introduction of .
+Here, STP and SFP denote True Positives and False Positives for “sure” predictions, while UTP and UFP are the corresponding values for the “unsure” category.
+For recall, we consider a prediction to be a UTP if it is partially correct (i.e., refinable).
+Unlike precision, we do not apply the weighting parameter to the denominator of the recall calculation, as recall measures the proportion of gold labels retrieved, which is independent of the introduction of .
 Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1].
@@
 2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
-The Weighted-F1 metric structurally aligns with our four-tiered alignment objective (Equation (3)) through a parametric formulation that encodes hierarchical priorities.
-Its design ensures: (1) Strict enforcement for sure predictions: Unweighted sure terms prioritize high-confidence correctness (STP) while imposing maximal penalties for sure errors (SFP); (2) Controlled utility for unsure predictions: -scaled unsure terms modulate low-confidence outputs—attenuated UFP penalties tolerate unsure errors, whereas partial UTP rewards incentivize refinable correct predictions.
+The Weighted-F1 metric aligns with our four-tiered objective (Equation (3)) by using a parametric formulation that encodes hierarchical priorities, which ensures: (1) Strict enforcement for sure predictions: Unweighted terms prioritize high-confidence correctness (STP) and apply maximal penalties for errors (SFP).
+(2) Controlled utility for unsure predictions: The -scaled terms modulate low-confidence outputs, with attenuated UFP penalties tolerating unsure errors, and partial UTP rewards incentivizing refinable correct predictions.
 This dual-parameter mechanism enables scenario-specific adaptation: Weighted-Precision regulates error tolerance via (higher tightens unsure-error constraints) and Weighted-Recall controls reference-value incentives via (higher promotes helpfulness through unsure-correct outputs).
 The composite metric thus operationalizes the reliability-helpfulness trade-off, permitting calibration from conservative (, ) to exploratory (, ) strategies.
-We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all ”unsure” predictions.
-(2) = 0, = 1: Errors in the ”unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement.
+We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all “unsure” predictions.
+(2) = 0, = 1: Errors in the “unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement.
 (3) = 1, = 0: Lacks significant interpretation.
@@
 2.2.3 Reliability Metrics
-We employ two specific Weighted-F1 scores as performance metrics. Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
-Quality-F1 Defined as , it assigns ”unsure” predictions half the weight of ”sure” ones, encouraging the model to select ”sure” when confident while minimizing penalties for ”unsure” outputs.
-Quality-F1 measures the model’s self-awareness ability and reliability.
-Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of ”unsure” predictions, reflecting the overall potential of the two-stage system.
+We use two specific Weighted-F1 scores: Quality-F1 as the primary metric for assessing model reliability, and Optimal-F1 as a supplementary reference.
+Quality-F1: Defined as , this metric assigns half the weight to “unsure” predictions compared to “sure” ones.
+This encourages the model to provide high-confidence “sure” outputs while mitigating penalties for “unsure” ones.
+Quality-F1 directly measures the model’s self-awareness and reliability.
+Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of “unsure” predictions, reflecting the overall potential of the system.
 3 Method
 3.1 Explicit Knowledge Boundary Modeling
-We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the ”unsure” predictions.
+We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve model reliability and self-awareness by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the “unsure” predictions.
 As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful.
+We provide a complete example of the two-stage execution flow and model outputs in Appendix B.
 Fast Prediction with Confidence Labeling
-In the first stage, the model performs the primary task while simultaneously assigning a confidence label (”sure” or ”unsure”) to each prediction.
-Predictions labeled as ”sure” are directly accepted as final outputs, ensuring high precision.
+In the first stage, the model performs the primary task and simultaneously assigns a confidence label (“sure” or “unsure”) to each prediction.
+“Sure” predictions are directly accepted as final outputs, ensuring high precision.
 Slow Refinement for Unsure Prediction
-In the second stage, ”unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
+In the second stage, “unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
 3.2 Reliability Alignment
-To align the model’s behavior with our proposed alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025).
-We apply Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for alignment, with the data construction process illustrated in Figure 2(2) and two detailed examples provided in Appendix §B.
+To align the model’s behavior with our alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025), including Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO).
+The data construction process is shown in Figure 2(2), with two detailed examples provided in Appendix §C.
 SFT Data Construction
-We assume that the model’s task-specific knowledge boundary can be approximated by repeatedly sampling on the training data.
-Accordingly, our SFT data construction approach is straightforward: we conduct multiple sampling for each training sample.
-Predictions consistently deemed accurate are labeled ”sure,” while those with errors or omissions are marked ”unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance.
-That is, when the model produces an incorrect prediction, we simply assign a confidence label of ”unsure” without rectifying the error.
+We assume that a model’s task-specific knowledge boundary can be approximated by repeated sampling on training data.
+Accordingly, our SFT data construction is straightforward: for each training sample, we perform multiple samplings.
+Predictions consistently deemed accurate are labeled “sure,” while those with errors or omissions are marked “unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance.
+That is, when the model produces an incorrect prediction, we simply assign a confidence label of “unsure” without rectifying the error.
 The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
-The number of sampling rounds, denoted as i, is a hyperparameter.
-A larger i leads to a more conservative model, as the proportion of ”unsure” labels increases with greater i in the training corpus.
+The number of sampling rounds i is a hyperparameter.
+A larger i results in a more conservative model, as the proportion of “unsure” labels in the training corpus increases.
 In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@
 We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction.
-This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §C.
+This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §D.
 4 Experiments
 We conduct experiments focusing on three core research questions:
-RQ1 Does our reliability training pipeline outperform current methods in enhancing self-awareness?
-RQ2 Does our EKBM framework improve the overall performance of LLMs on complex tasks?
+RQ1 Does our training pipeline outperform current methods in enhancing self-awareness?
+RQ2 Does EKBM framework improve the overall performance of LLMs on complex tasks?
 RQ3 Does our approach scale robustly across foundation models with varying capabilities?
@@
 Training details could be found in Appendix §A.
-To evaluate the reliability of our proposed models, we compare them against several baselines, including three prompt-based and three uncertainty-based approaches (see Appendix §D).
-Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models for comparison.
-By default, we utilize only 1,000 samples for Reliability-SFT training and 2,000 for DPO training—significantly less than traditional models.
+To evaluate the reliability of our proposed models, we compare them against six baselines: three prompt-based and three uncertainty-based approaches (see Appendix §E).
+Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models.
+By default, we use only 1,000 samples for Reliability-SFT training and 2,000 for DPO, which is significantly less than traditional models.
 Prompt-based
-methods encompass three approaches: Direct, Verbose, and Self-Reflection (SR).
-The Direct approach generates predictions without confidence estimation, whereas Verbose simultaneously produces confidence labels.
+methods encompass three approaches: Direct, Verbose, and Self-Reflection (denoted as SR, Ji et al.
+(2023a)).
+The Direct approach generates predictions without confidence estimation, while Verbose produces both predictions and confidence labels simultaneously.
 SR assesses each prediction through self-evaluative reflection.
 Uncertainty-based
-methods include Token Probability (Prob), Self-Consistency (SC), and P(True) (Kadavath et al., 2022), each employing distinct techniques to classify outputs as ”sure” or ”unsure”.
-Prob calculates average token probabilities, SC measures the frequency of predictions across multiple samples, and P(True) determines the probability of the ”True” token during model self-evaluation.
-Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Dubey et al.
-(2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B Yang et al.
-(2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
+methods include Token Probability (Prob, Manakul et al.
+(2023)), Self-Consistency (SC, Chen & Mueller (2024)), and P(True) (Kadavath et al., 2022).
+These methods use different techniques to classify outputs as “sure” or “unsure”.
+Prob calculates average token probabilities; SC measures the frequency of predictions across multiple samples; and P(True) determines the probability of the “True” token during model self-evaluation.
+Method Type Method MultiWOZ-2.4 BiTOD SGD Opti.
+F1 Qual.
+F1 Opti.
+F1 Qual.
+F1 Opti.
+F1 Qual.
+F1 LLAMA3-8B (Dubey et al., 2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B (Yang et al., 2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
 4.1.2 Dataset
@@
 For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, demonstrating a substantial improvement in self-awareness.
-Our approach consistently achieves superior Quality-F1 scores, indicating improved effectiveness in distinguishing predictions as ”sure” or ”unsure” based on confidence levels.
-Notably, a high indicates reliable ”sure” predictions, fostering user trust with minimal need for verification.
-Simultaneously, a high ensures that the ”unsure” category encompasses a broad spectrum of potentially correct outputs, forming a solid basis for further refinement.
-Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance.
-In other words, our system optimally balances precision and recall, delivering both immediate reliability and greater potential for post-refinement accuracy.
-In our approach, methods incorporating DPO generally outperform SFT-only methods.
-Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores.
-This suggests that Joint Training models adopt a more conservative stance, deferring to subsequent refinement, whereas Post Training models exhibit greater confidence, producing predictions of generally higher quality.
-This discrepancy stems from the differing sampling distributions of DPO data: Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of ”unsure” predictions and resulting in more cautious model behavior.
-Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority.
-However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as ”unsure.” This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score.
-In extreme cases, a model could assign ”unsure” to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
+Our approach consistently achieves superior Quality-F1 scores, demonstrating improved ability to distinguish predictions as “sure” or “unsure” based on confidence.
+A high indicates reliable “sure” predictions, fostering user trust with minimal need for verification.
+Simultaneously, a high ensures the “unsure” category contains a broad range of potentially correct outputs, which forms a solid basis for refinement.
+Together, these metrics lead to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance.
+In essence, our system optimally balances precision and recall, providing both immediate reliability and a strong potential for post-refinement accuracy.
+In our approach, DPO-based methods generally outperform SFT-only methods.
+Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, suggesting a more conservative stance that defers to subsequent refinement.
+In contrast, those undergoing Post Training yield superior Quality-F1 scores, indicating greater confidence and higher-quality predictions.
+This discrepancy arises because Joint Training samples from a pre-SFT distribution, which has poorer performance.
+This results in a dataset with a higher proportion of “unsure” predictions, leading to more cautious model behavior.
+Notably, some baselines, such as Self-Consistency, show high , , and Optimal-F1 scores, suggesting initial superiority.
+However, this often indicates an overly conservative model that frequently labels predictions as “unsure.” This leads to a suboptimal balance between prediction types, as evidenced by a significantly lower Quality-F1 score.
+In an extreme case, a model could label all predictions as “unsure”, maximizing Optimal-F1 but shifting the entire burden to the refinement stage.
+This approach sacrifices immediate usability and fails to balance reliability and helpfulness.
 4.3 Refinement for Unsure Prediction
-We trained an automatic refinement model for each dataset as described in Section 3.3.
-We refine the predictions labeled as ”unsure” from the initial stage and merge the refined outputs with the ”sure” predictions to generate the final results.
-The overall evaluation results are presented in Table 2.
-For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
-Method Type | Method | MultiWOZ-2.4 | BiTOD | SGD | |||
-LLAMA3-8B | |||||||
-Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 |
-Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |
-SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |
-Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 |
-SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |
-P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |
-Reliability (Ours) | SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 |
-+ DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |
-+ DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |
-Qwen2.5-7B | |||||||
-Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 |
-Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |
-SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |
-Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 |
-SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |
-P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |
-Reliability (Ours) | SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 |
-+ DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |
-+ DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 |
-It is important to note that, referring to §4.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models.
+As described in Section §3.3, we trained an automatic refinement model for each dataset.
+We refined the “unsure” predictions from the initial stage and merged them with the “sure” predictions to generate the final results.
+Table 2 shows the overall evaluation.
+For RQ2, our method achieves significantly better performance after refinement, with substantial improvements in Slot-F1 and JGA.
+MultiWOZ-2.4 | BiTOD | SGD | |||||
+| Slot-F1 | JGA | Slot-F1 | JGA | Slot-F1 | JGA | ||||
+| LLAMA3-8B | |||||||||
+| Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 | ||
+| Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |||
+| SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |||
+| Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 | ||
+| SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |||
+| P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |||
+SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 | ||
+| + DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |||
+| + DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |||
+| Qwen2.5-7B | |||||||||
+| Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 | ||
+| Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |||
+| SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |||
+| Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 | ||
+| SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |||
+| P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |||
+SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 | ||
+| + DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |||
+| + DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 | |||
+Referring to §4.2, despite their higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement compared to Post Training models.
 This discrepancy stems from the limitations of the refinement model.
-Since the refinement model is not ideal, excessive ”unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains.
-This highlights the importance of a balanced ”sure” and ”unsure” classification.
-An ideal model should minimize unnecessary ”unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
+Since the refinement model is not perfect, an excessive number of “unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains.
+This highlights the importance of a balanced “sure” and “unsure” classification.
+An ideal model should minimize unnecessary “unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
 4.4 Scalability Analysis
-To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities.
-We consider the original LLAMA3 8B model a lower-performance baseline (denoted as Low) and applied supervised fine-tuning (SFT) on each dataset to enhance its DST capabilities.
-We train two foundation models using 1,000 samples (Medium) and 10,000 samples (High), then continually conduct reliability training and evaluate the performance of the obtained reliable LLM against baseline algorithms, with results shown in Figure 3.
-For RQ3, our methods consistently outperform baselines across different foundation models after refinement, demonstrating strong scalability—particularly in real-world scenarios where models are often domain-specific tuned and possess high task capabilities.
-Detailed results in Appendix §E further support these findings.
-Additionally, as illustrated in Appendix Table 5, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves.
-This trend stems from the offline nature of our DPO algorithm—higher-performing foundation models exacerbate the limitations of the Joint Training data distribution compared to Post Training, thereby affecting overall performance.
-Notably, in the SGD dataset, the benefits of DPO Joint Training diminish as the foundation model’s performance increases, even underperforming the SFT-only approach.
+To evaluate scalability, we experimented with foundation models of varying capabilities.
+We consider the original LLAMA3 8B as a low-performance baseline (denoted as Low), then fine-tuned two models with 1,000 (Medium) and 10,000 (High) samples to enhance their DST capabilities.
+We then continually conducted reliability training and evaluated the performance of the obtained reliable LLMs against baseline methods.
+As shown in Figure 3, our methods consistently outperform baselines after refinement, demonstrating strong scalability.
+This is particularly relevant for real-world scenarios where models are often domain-specifically tuned and possess high task capabilities.
+Detailed results in Appendix §F further support these findings.
+As illustrated in Table 7 in Appendix §F, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves.
+This is because our offline DPO algorithm’s limitations are exacerbated by higher-performing foundation models, negatively affecting the Joint Training data distribution and overall performance.
+Notably, in the SGD dataset, the benefits of DPO Joint Training diminish with increasing foundation model performance, even falling behind the SFT-only approach.
 4.5 DPO Preference Strategies
-We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1.
-Weighted-F1 incorporates two adjustable parameters: influences Weighted Precision, while affects Weighted Recall.
-Using the Reliability-SFT trained LLAMA3 8B model as backbone, we further generated DPO preference data with varying values and conducted consecutive Post Training.
-Test results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance with and , reaching near-best results at a relatively low cost.
-Analysis indicates that increasing heightens sensitivity to errors in ”unsure” predictions, reducing tolerance and yielding fewer ”unsure” outputs.
-Conversely, increasing amplifies rewards for successfully recalling ”unsure” predictions, encouraging the model to produce more ”unsure” predictions and adopt a more conservative stance.
-Practically, the optimal alpha settings depend on the scenario and refinement model performance.
-If the refinement model underperforms, excessive costs for marginal gains may not be justified.
+We constructed our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1.
+Weighted-F1 incorporates two adjustable parameters: for Weighted Precision and for Weighted Recall.
+Using the Reliability-SFT trained LLAMA3 8B model, we generated DPO preference data with varying values and performed subsequent Post Training.
+Results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance at and , reaching near-best results at a relatively low cost.
+Analysis indicates that increasing heightens sensitivity to errors in “unsure” predictions, reducing tolerance and the number of “unsure” outputs.
+Conversely, increasing amplifies rewards for successfully recalling “unsure” predictions, encouraging a more conservative stance.
+The optimal alpha settings depend on the specific scenario and refinement model performance; if the refinement model underperforms, the cost of excessive “unsure” predictions may not be justified.
 The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
@@
 4.6.1 Prediction Types Analysis
-We further analyzed the model’s behavior after Reliability Training to evaluate its ability to make reliable confidence judgments in its outputs.
-Specifically, we examined the distribution of output categories and compared them to baseline models.
-These categories include five cases: correct and incorrect predictions classified as ”sure” or ”unsure”, along with missing predictions.
+To evaluate our model’s ability to make reliable confidence judgments, we further analyzed the distribution of its output categories and compared them to baselines.
+We examined five output categories: correct and incorrect predictions classified as “sure” or “unsure”, along with missing predictions.
 The results are presented in Figure 5.
-Our findings highlight two key improvements: (1) Enhanced reliability—The proportion of incorrect predictions within the ”sure” category has been notably reduced, with most errors now concentrated in the ”unsure” category.
-This indicates a stronger self-assessment capability, allowing the model to better identify potentially erroneous predictions compared to baselines.
-Additionally, across all scenarios, the ”unsure” category contains fewer correct predictions than incorrect ones, demonstrating the model’s effectiveness in placing accurate outputs in the ”sure” category, a distinction baseline models struggle to achieve.
-(2) Improved recall—Our approach significantly reduces missing predictions, capturing previously omitted relevant information within the ”unsure” category.
-While some may still contain errors, this shift enhances overall system helpfulness and provides a stronger foundation for further refinement.
+Our findings highlight two key improvements: (1) Enhanced reliability: We observed a notable reduction in incorrect “sure” predictions, with most errors concentrated in the “unsure” category.
+This indicates a stronger self-assessment capability, allowing the model to better identify potential errors than baselines.
+Additionally, the “unsure” category consistently contains fewer correct predictions than incorrect ones, a distinction baseline models struggle to achieve.
+(2) Improved recall: Our approach significantly reduces missing predictions by capturing previously omitted relevant information within the “unsure” category.
+While some of these may still contain errors, this shift enhances overall helpfulness and provides a stronger foundation for refinement.
 4.6.2 Cost Analysis
-In this work, ”Cost” refers to the additional overhead from the refinement process.
-We define the cost of refining an ”unsure” prediction as a constant value of 1, equating Cost to the number of ”unsure” predictions.
-For better comparison, we express Cost as the proportion of ”unsure” predictions.
-We compare the performance and Cost of various baselines, as shown in Figure 6.
-Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix §F.
-Our methods demonstrate significant performance improvements while maintaining a relatively low Cost.
-Furthermore, as illustrated in Appendix Table 7, Cost decreases significantly with improved foundation model performance.
-This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
+In this work, Cost refers to the additional overhead from the refinement process.
+We define the cost of refining an “unsure” prediction as a constant value of 1, equating Cost to the number of “unsure” predictions.
+For better comparison, we express Cost as the proportion of “unsure” predictions.
+We compare the performance and Cost of various baselines in Figure 6.
+Results are shown for the MultiWOZ-2.4 dataset, with complete findings in Appendix §G.
+Our methods show significant performance improvements at a relatively low Cost.
+Additionally, as shown in Appendix Table 9, Cost decreases significantly as the foundation model’s performance improves, which is expected as a stronger model produces fewer uncertain predictions.
 4.6.3 Refinement Model Comparison
-We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 7.
-In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series (DeepSeek-AI et al., 2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities.
-The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o.
-Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
+We used various refinement models to enhance predictions from the LLAMA3 Reliability-SFT+DPO Post model.
+As shown in Table 7, we compared our trained models against GPT-4o and the DeepSeek series (DeepSeek-AI et al., 2025) (8B and 70B), all known for strong reasoning.
+Results show significant performance differences, with DeepSeek 8B underperforming while the 70B model nearly matches GPT-4o.
+Our task-specific fine-tuning with CoT substantially improves refinement accuracy.
+4.7 Extension to Mathematical Problems
+To validate the generalization of our method to different tasks, we conducted supplementary experiments on the widely used GSM8K mathematics dataset (Cobbe et al., 2021) using the Qwen2.5-7B-Instruct model with 1,000 training samples.
+We evaluated the model under two test settings: (a) with Chain-of-Thought (CoT) reasoning (Wei et al., 2023) and (b) by directly predicting the final answer.
+We tested both settings as LLMs generally perform differently on each.
+For the refinement model, we used the DeepSeek-R1-Distill-Llama-70B model, which has strong mathematical reasoning capabilities.
+We adapted the F1-based evaluation metric from Section §2.2 to an accuracy-based definition:
+As shown in Table 3, our method consistently outperformed both prompt-based and uncertainty-based baselines in improving the model’s self-assessment ability, achieving higher Quality-Accuracy.
+Furthermore, after applying refinement with the same refinement model, our approach yielded a higher final accuracy at a reasonable computational cost.
+| Method Type | Method | Quality Accuracy | Final Accuracy | ||
+| Prompt | Direct | 86.81 | 86.81 | ||
+| Verbose | 90.8 | 90.98 | |||
+| Uncertainty | Prob | 87.69 | 88.55 | ||
+| SC | 89.41 | 91.48 | |||
+SFT | 89.69 | 91.05 | ||
+| + DPO Joint | 90.14 | 93.10 | |||
+| + DPO Post | 92.73 | 95.30 |
+| Method Type | Method | Quality Accuracy | Final Accuracy | ||
+| Prompt | Direct | 19.94 | 19.94 | ||
+| Verbose | 21.44 | 24.34 | |||
+| Uncertainty | Prob | 24.49 | 76.50 | ||
+| SC | 24.37 | 70.05 | |||
+SFT | 27.31 | 77.94 | ||
+| + DPO Joint | 26.16 | 78.24 | |||
+| + DPO Post | 27.83 | 79.83 |
 5 Related Work
@@
 Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a).
-While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024).
-To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023), albeit with increased computational overhead.
+While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024; Pan et al., 2025; Shan et al., 2025).
+To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023b), albeit with increased computational overhead.
 More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024).
-While effective in mitigating misinformation, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
+While effective, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
 6 Conclusion
@@
 Extensive evaluations demonstrate that our method significantly improves model self-awareness and performance across complex tasks, establishing a new paradigm for enhancing the reliability of language models in error-sensitive applications.
+Acknowledgments
+This work is funded by the China NSFC Projects (62120106006, 92370206, and U23B2057) and Shanghai Municipal Science and Technology Projects (2021SHZDZX0102 and 25X010202846).
 Appendix A Training Details
@@
 All models are trained for 1 epoch using NVIDIA A800 GPUs.
-Appendix B Training Data Construction
+Appendix B Execution Example for Two Stages
+We provide an example sample that includes the complete execution flow and model outputs for the two stages as shown in Table 4.
+| Input | ||||||||
+| Stage One: Reliable LLM Predictions | ||||||||
+| Stage Two: Refinement for “Unsure” Predictions: | ||||||||
+| Final Refined Predictions: | ||||||||
+Appendix C Training Data Construction
 We provide detailed examples to illustrate our training data construction process, including model inputs and the expected outputs the model is intended to learn.
 The data construction for SFT is outlined in Table 5, while that for DPO is detailed in Table 6.
-Input |
-Dialog history: USER: Is there a modern European restaurant in the north?
-SYSTEM: I found city stop restaurant.
-They are located at Cambridge City Football Club Milton Road Chesterton.
-USER: That sounds lovely!
-SYSTEM: Would you like me to help you with reserving a table?
-USER: Maybe.
-Is that restaurant in the moderate price range?
-SYSTEM: No, it is in the expensive price range.
-Is this okay, or would you like to try another area?
-USER: How about a restaurant that serves Indian food?
-SYSTEM: There are several restaurants what part of town would you like it?
-USER: A restaurant in the north, please.
-Reliable dialog state: |
-Gold Labels |
-{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
-LLM Original Predictions |
-{’restaurant’: {’area’: ’north’, ’food’: ’european’, ’name’: ’Cambridge City Football Club Milton Road Chesterton’}} |
-Constructed SFT Sample Output |
-{’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’European’, ’confidence’: ’unsure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}, ’name’: {’value’: ’Cambridge City Football Club Milton Road Chesterton’, ’confidence’: ’unsure’}}} |
-Explanation |
-In the LLM’s original predictions, the value for area is correct, food is incorrect, name is extraneous, and pricerange is missing.
-Consequently, in the constructed SFT sample output, area is labeled as ”sure,” while the other three are marked as ”unsure,” with erroneous information left uncorrected.
-|
-Input |
-Dialog history: USER: Is there a modern European restaurant in the north?
-SYSTEM: I found city stop restaurant.
-They are located at Cambridge City Football Club Milton Road Chesterton.
-USER: That sounds lovely!
-SYSTEM: Would you like me to help you with reserving a table?
-USER: Maybe.
-Is that restaurant in the moderate price range?
-SYSTEM: No, it is in the expensive price range.
-Is this okay, or would you like to try another area?
-USER: How about a restaurant that serves Indian food?
-SYSTEM: There are several restaurants what part of town would you like it?
-USER: A restaurant in the north, please.
-Reliable dialog state: |
-Gold Labels |
-{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
-Multi-Sampling With the Foundation LLM |
-Constructed DPO Sample Output Pairs |
-Chosen {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}}} Rejected {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}}} |
-Explanation |
-Compared to the rejected sample, the chosen sample exhibits a higher recall for a correct pricerange within the ”unsure” category, aligning more closely with our alignment objectives.
-This is reflected in a higher Weighted-F1 score, justifying its selection over the rejected sample.
-|
-Appendix C Refinement Model
+| Input | |||||||||||
+| Gold Labels | |||||||||||
+| LLM Original Predictions | |||||||||||
+| Constructed SFT Sample Output | |||||||||||
+| Explanation | |||||||||||
+| Input | |||||||||||
+| Gold Labels | |||||||||||
+| Multi-Sampling With the Foundation LLM | |||||||||||
+| Constructed DPO Sample Output Pairs | |||||||||||
+| Explanation | |||||||||||
+Appendix D Refinement Model
 We trained an automated refinement model for each dataset.
@@
 For each dataset, we sampled 5,000 examples from the model’s unsure predictions as well as from randomly selected data in the training set, ensuring a balanced distribution of data across three actions.
-Appendix D Baselines Details
+Appendix E Baselines Details
 Here we detail the specifics of the baselines.
@@
 Uncertainty-based
-Token Probability (Prob): Computes average token-level probabilities, classifying outputs as ”sure” or ”unsure” based on a threshold , with probabilities normalized and set to 0.8.
-Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as ”sure,” while others are marked as ”unsure.” We conduct ten samples () and set to 50%.
-P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding ”True” than ”False” as ”sure,” otherwise as ”unsure.”
-Appendix E Scalability Analysis
+Token Probability (Prob): Computes average token-level probabilities, classifying outputs as “sure” or “unsure” based on a threshold , with probabilities normalized and set to 0.8.
+Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as “sure,” while others are marked as “unsure.” We conduct ten samples () and set to 50%.
+P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding “True” than “False” as “sure,” otherwise as “unsure.”
+Appendix F Scalability Analysis
 The detailed results are presented in Table 7, demonstrating that our method maintains superiority across various capability baselines, achieving more reliable outcomes (higher , , and Quality-F1) and enhanced overall performance post-refinement (higher JGA).
-Method Type Method MultiWOZ-2.4 BiTOD SGD 1000Samples-Trained LLAMA3 Based Prompt Direct 91.07 86.28 87.84 42.73 93.01 92.78 92.82 70.91 82.28 85.10 82.77 38.13 Uncertainty Prob 91.84 86.52 86.84 45.43 94.93 93.55 93.66 72.73 84.45 85.24 82.80 39.48 SC 92.68 92.37 84.31 43.18 95.70 95.12 92.05 70.37 86.61 92.30 78.71 28.62 Reliability (Ours) SFT 95.42 92.11 89.86 55.62 94.97 94.74 93.80 73.64 85.16 85.14 81.06 39.23 + DPO Joint 95.11 90.68 89.05 58.94 96.06 95.60 95.11 75.36 81.66 82.82 79.69 37.62 + DPO Post 96.33 92.15 89.91 59.12 96.22 95.54 95.43 77.70 86.67 85.44 83.27 41.48 10000Samples-Trained LLAMA3 Based Prompt Direct 94.51 93.30 93.63 63.19 95.35 95.14 94.34 80.76 83.24 83.75 83.11 43.96 Uncertainty Prob 95.77 94.06 92.95 67.39 95.95 95.53 95.23 81.26 84.37 84.61 83.08 43.52 SC 96.24 96.40 92.13 67.61 97.83 97.81 93.42 78.04 88.13 92.42 79.26 38.20 Reliability (Ours) SFT 96.98 95.86 94.40 70.85 96.44 96.20 96.03 80.54 85.30 83.92 83.57 45.11 + DPO Joint 96.98 95.61 94.71 70.84 97.15 96.65 95.96 82.35 82.31 82.36 80.75 40.48 + DPO Post 97.71 96.32 94.80 71.51 96.57 96.35 96.59 82.69 86.29 85.53 84.21 45.43
-Appendix F Refinement Cost
-The detailed overall results of cost are presented in Table 7 and 7.
-The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 8(a) and Figure 8(b), respectively.
+Method Type Method MultiWOZ-2.4 BiTOD SGD Qual.
+F1 JGA Qual.
+F1 JGA Qual.
+F1 JGA 1000Samples-Trained LLAMA3 Based Prompt Direct 91.07 86.28 87.84 42.73 93.01 92.78 92.82 70.91 82.28 85.10 82.77 38.13 Uncertainty Prob 91.84 86.52 86.84 45.43 94.93 93.55 93.66 72.73 84.45 85.24 82.80 39.48 SC 92.68 92.37 84.31 43.18 95.70 95.12 92.05 70.37 86.61 92.30 78.71 28.62 Reliability (Ours) SFT 95.42 92.11 89.86 55.62 94.97 94.74 93.80 73.64 85.16 85.14 81.06 39.23 + DPO Joint 95.11 90.68 89.05 58.94 96.06 95.60 95.11 75.36 81.66 82.82 79.69 37.62 + DPO Post 96.33 92.15 89.91 59.12 96.22 95.54 95.43 77.70 86.67 85.44 83.27 41.48 10000Samples-Trained LLAMA3 Based Prompt Direct 94.51 93.30 93.63 63.19 95.35 95.14 94.34 80.76 83.24 83.75 83.11 43.96 Uncertainty Prob 95.77 94.06 92.95 67.39 95.95 95.53 95.23 81.26 84.37 84.61 83.08 43.52 SC 96.24 96.40 92.13 67.61 97.83 97.81 93.42 78.04 88.13 92.42 79.26 38.20 Reliability (Ours) SFT 96.98 95.86 94.40 70.85 96.44 96.20 96.03 80.54 85.30 83.92 83.57 45.11 + DPO Joint 96.98 95.61 94.71 70.84 97.15 96.65 95.96 82.35 82.31 82.36 80.75 40.48 + DPO Post 97.71 96.32 94.80 71.51 96.57 96.35 96.59 82.69 86.29 85.53 84.21 45.43
+Appendix G Refinement Cost
+The detailed overall results of cost are presented in Table 9 and 9.
+The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 8(a) and Figure 8(b), respectively, where cost refers to the proportion of unsure predictions.
 Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA Qwen LLAMA Qwen LLAMA Qwen Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Verbose 87.28 30.62 80.69 38.05 90.30 74.07 SR 55.43 28.07 95.50 8.55 84.77 15.07 Uncertainty Prob 36.47 40.65 50.46 52.91 56.68 53.01 SC 69.26 58.91 58.05 62.59 96.80 92.71 P(True) 14.63 10.22 13.35 1.82 1.52 1.71 Reliability (Ours) SFT 32.67 30.98 20.03 21.78 51.24 55.96 + DPO Joint 31.96 31.98 21.19 23.63 57.72 45.29 + DPO Post 27.69 28.38 10.54 18.14 40.68 53.10
 Method Type Method MultiWOZ-2.4 BiTOD SGD Medium High Medium High Medium High Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Uncertainty Prob 6.55 5.74 3.29 1.37 8.14 5.53 SC 28.93 12.17 6.25 2.12 41.68 41.42 Reliability (Ours) SFT 13.52 5.12 2.16 0.66 12.60 1.87 + DPO Joint 10.70 3.94 1.36 1.04 11.66 2.67 + DPO Post 13.32 8.59 3.44 1.46 19.01 4.19
-Refinement Examples |
-Give you a dialog history between USER and SYSTEM, and a slot-value pair extracted from the dialog, I want you to analyze on the dialog history and review/refine the given dialog state slot-value pair to make them more accurate and reliable.
-## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Question Dialog history: USER: Hello.
-I need train to London liverpool Street.
-SYSTEM: Where are you departing from and do you have a time preference?
-USER: Yes, I’d like to leave Cambridge sometime after 7:00.
-SYSTEM: I have one that leaves at 7:59 and 4 more that depart every two hours after.
-USER: The 7:59 will be fine.
-Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: ”leaveat” is one of the slots in the train domain and the value is incorrectly specified as ”7:59” (formation error).
-Precise time should strictly follow the ”HH:MM” format.
-step2: semantic aspect: ”leaveat” is not a ”name”-related slot, so the value must be extracted from the user’s direct intent.
-Although the system provides a train that departs at 7:59, the user only mentions that he’d like to leave Cambridge after 7:00.
-Considering the formation, the correct value should be 07:00.
-step3: domain train correct; slot leaveat correct; value 7:59 incorrect.
-step4: the predicted slot-value pair is incorrect in both structure and semantic, and the value should be corrected.
-Refine action: correct Refined slot-value pair: train: leaveat = 07:00 |
-GPT4o Prompt For Refinement Thinking Steps Generation |
-Give you a dialog history between USER and SYSTEM, a predicted slot-value pair and a gold refinement, I want you to analyze on the dialog history, refer to the examples I’ve provided and the golden refinement, and then continue to provide the thinking steps following the predefined principles.
-Firstly, we use json dict to describe the slots and their corresponding value space.
-Then, we will specify the requirements you need to comply and provide some examples.
-Last, we will present you with the dialog history, the predicted dialog state slot-value pair and the gold refinement for you to perform continuation.
-## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Meta Requirements 1.
-Analyze the dialog history and the predicted slot-value pair carefully, correct and refine the pair according to the following Detailed Refinement Principles.
-2.
-For the given slot-value pair, you should only take one of the actions of ”reserve”, ”remove” or ”correct” according to the following principles.
-a.
-reserve: if the slot is relevant to the dialog history and the value is correctly specified, you should keep the pair.
-b.
-remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print ”none” indicating that the pair has been removed.
-c.
-correct: if the slot is relevant to the dialog history and the value is incorrectly specified, you should extract the correct value and replace it.
-3.
-You need to think and reason in the same way as the Examples.
-Please refer to the Examples for formatting requirement.
-Make sure the output is all lowercase.
-Only focus on the given slot-value pair and do not consider any other slot.
-only provide the reasonable thinking steps, do not provide any extra prefixes or suffixes.
-## Detailed Refinement Principles ## Examples ## Question Dialog history: USER: … SYSTEM: … Predicted slot-value pair: train: leaveat = 7:59 Refinement: Refine action: correct Refined slot-value pair: train: leaveat = 07:00 Thinking steps: |
-Appendix G LLM Instructions
+| Refinement Examples | |||||||||||||||||||||||||
+| GPT4o Prompt For Refinement Thinking Steps Generation | |||||||||||||||||||||||||||||||||
+Appendix H LLM Instructions
 The instructions used in this work are presented in this chapter, taking the MultiWOZ-2.4 dataset as an example. The instructions for the BiTOD dataset are similar.
-G.1 Inference Instruction
-The prompts used with the original untrained LLAMA3 model are shown in Tables 10 and 11.
+H.1 Inference Instruction
+The prompts used with the original untrained LLAMA3 model are shown in Tables 12 and 13.
 Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
-Direct Inference Instruction |
-Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state.
-Firstly, we use json dict to describe the slots and their corresponding value space in each domain.
-Then, we will specify the requirements you need to comply.
-Last, we demonstrate some use cases.
-## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1.
-Analyze the dialog history carefully and fill the relevant domains and slots.
-2.
-For slots with a specified value range, responses must fall within the provided range.
-For slots without specified value range, the answer must be extracted from the history.
-Set value as ”dontcare” if user doesn’t have a preference.
-3.
-You only need to consider the domains and slots that are relevant to the conversation history.
-Do not include those irrelevant in your response and avoid presenting empty domains or slots.
-4.
-Your answer should also be in one-line jsonl format and make sure the output is all lowercase.
-Do not provide any extra prefixes or suffixes or any explanations.
-## Output Dialog State Example {”hotel”: {”area”: ”centre”, ”name”: ”alexander bed and breakfast”, ”parking”: ”yes”, ”type”: ”guesthouse”}, ”attraction”: {”name”: ”kambar”}} ## Examples shots… |
-Reliability Inference Instruction |
-Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state.
-Firstly, we use json dict to describe the slots and their corresponding value space in each domain.
-Then, we will specify the requirements you need to comply.
-Last, we demonstrate some use cases.
-## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1.
-Analyze the dialog history carefully and fill the relevant domains and slots.
-2.
-For slots with a specified value range, responses must fall within the provided range.
-For slots without specified value range, the answer must be extracted from the history.
-Set value as ”dontcare” if user doesn’t have a preference.
-3.
-You should try to cover as many domians and slot-value pairs relevant to the conversation history as possible.
-Mark each slot with either ”sure” or ”unsure” according to your confidence in it.
-For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with ”sure”.
-Otherwise, you should tag the slot-value pair with ”unsure”.
-4.
-Principles: a.
-Goal one: Achieving as close to 100% accuracy as possible in those ”sure” slot-value pairs.
-b.
-Goal two: The ”sure” part along with the ”unsure” part should cover all possible slots involved in the dialog history as much as possible.
-c.
-Heavy penalization: Providing incorrect slot-value pairs in the ”sure” part.
-d.
-Heavy penalization: Missing slot-value pairs that should be extracted.
-e.
-Light penalization: Providing incorrect or redundant slot-value pairs in the ”unsure” part.
-5.
-Your answer should also be in one-line jsonl format and make sure the output is all lowercase.
-Do not provide any extra prefixes or suffixes or any explanations.
-Please refer to the Output Dialog State Example for formatting requirement.
-## Output Dialog State Example {”hotel”: {”area”: {”value”: ”centre”, ”confidence”: ”sure”}, ”name”: {”value”: ”alexander bed and breakfast”, ”confidence”: ”sure”}, ”parking”: {”value”: ”yes”, ”confidence”: ”unsure”}, ”type”: {”value”: ”guesthouse”, ”confidence”: ”sure”}}, ”attraction”: {”name”: {”value”: ”kambar”, ”confidence”: ”unsure”}}} ## Examples shots… |
-G.2 Training Instruction
-The prompts used with the trained LLAMA3 model are shown in Tables 12 and 13.
+| Direct Inference Instruction | |||||||||||||||||||||||
+| Reliability Inference Instruction | |||||||||||||||||||||||||||||
+H.2 Training Instruction
+The prompts used with the trained LLAMA3 model are shown in Tables 14 and 15.
 The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results.
 In contrast, the Reliability Training Instruction is designed to train models capable of explicitly outputting confidence labels alongside predictions.
-Direct Training Instruction |
-Generate the dialogue state based on the given dialogue context. |
-Reliability Training Instruction |
-Generate the dialogue state based on the given dialogue context.
-Ensure the results are as reliable as possible, the ’sure’ parts are as accurate as possible, and the overall coverage includes all relevant slots.
-|
+| Direct Training Instruction | |
+| Reliability Training Instruction | |